### v4

# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

Abstract.
We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of cohomologically calibrated metric connections with totally skew torsion on product manifolds. We show that the dimension of the off-diagonal curvature subspace of a connection $\nabla^{C}$ is bounded below by the sum of tensor ranks of the mixed Künneth components of its calibration class, minus the dimension of the space of harmonic parallel forms on the first factor. The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}\in H^{3}(M;\mathbb{R})$ and on the space of $g_{1}$-parallel harmonic $2$-forms on $M_{1}$; it is therefore independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves parallel $2$-forms, hence it is metric-invariant modulo parallel-form preserving deformations. This provides a computational criterion for geometric complexity and quantifies the interaction between topology and curvature, yielding a quantified version of “forced irreducibility” via the dimension of $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
Keywords: Connections with Torsion, Product Manifolds, Holonomy Algebra, De Rham Cohomology, Hodge Theory, Künneth Decomposition, Curvature Tensor
MSC Classification 2020: 53C05, 53C07, 53C29, 58A14, 57R19, 15A69
1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological quantity, the mixed tensor rank of the cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that it furnishes, for any metric connection $\nabla^{\mathcal{C}}$, a quantified lower bound for the dimension of the subspace of the holonomy algebra generated by off-diagonal curvature operators with respect to the splitting $V_{1}\oplus V_{2}$. The key point is that, although the definition of “off-diagonal” uses $g$, the lower bound is independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves the space of parallel 2-forms, and is explicitly corrected by the dimension of $g_{1}$-parallel harmonic 2-forms on the first factor. Each independent mixed Künneth component in $\omega]$ guarantees a linearly independent off-diagonal curvature direction, modulo the parallel subspace.
2. Cohomologically Calibrated Metric Connections with Skew Torsion
Throughout, $\nabla^{LC}$ denotes the Levi-Civita connection of $g$. We consider metric connections $\nabla^{C}$ with totally skew torsion $T$, i.e.
$\nabla^{C}g=0,$ (2.1)
$T(X,Y,Z):=g(T(X,Y),Z)\in\Omega^{3}(M),$ (2.2)
and write
$\nabla^{C}=\nabla^{LC}+K,$ (2.3)
$K_{XYZ}=\tfrac{1}{2}\,T_{XYZ}.$ (2.4)
Thus $T^{\flat}:=T$ is a 3-form. The curvature $R^{C}$ of $\nabla^{C}$ can be expressed in terms of $R^{LC}$, $\nabla^{LC}T$, and quadratic torsion terms; in abstract index notation,
$R^{C}_{XY}Z=R^{LC}_{XY}Z+\tfrac{1}{2}\big((\nabla^{LC}_{X}T)(Y,Z,\cdot)^{\sharp}-(\nabla^{LC}_{Y}T)(X,Z,\cdot)^{\sharp}\big)+\tfrac{1}{4}\,Q_{T}(X,Y)Z,$ (2.5)
where $Q_{T}$ is bilinear and quadratic in $T$ (see, e.g., [3] or standard torsion-curvature formulas). We shall use (2.5) only qualitatively: mixed components of $T$ produce mixed components of $R^{C}$ through both the $\nabla^{LC}T$ and $T\ast T$ terms.
Definition 2.1 (Metric cohomological calibration).
A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the cohomology class $\omega]$ is non trivial and defined via the harmonic projection of the torsion three-form, ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, so $\omega]:=[\omega_{h}]$), is called cohomologically calibrated. The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
$H^{3}(M)\cong\bigoplus_{p+q=3}H^{p}(M_{1})\otimes H^{q}(M_{2}).$ (2.6)
Since $M$ is compact and oriented, Hodge theory applies to $g$: each cohomology class has a unique harmonic representative. We denote by $\omega_{h}$ the harmonic representative of $\omega]$ with respect to $g$.
For a thorough treatment of Hodge theory and cohomology on complex manifolds, see [5].
3. Off-Diagonal Curvature and Holonomy Complexity
Let $P_{i}:T_{p}M\to V_{i}$ be the orthogonal projections of the product splitting.
Definition 3.1 (Off-diagonal curvature).
For a curvature tensor $R$, define its off-diagonal component by
$R_{\mathrm{off}}(X,Y)Z:=P_{1}R(X,Y)(P_{2}Z)+P_{2}R(X,Y)(P_{1}Z).$ (3.1)
Definition 3.2 (Off-diagonal holonomy subspace).
The off-diagonal holonomy subspace at $p\in M$ is
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\operatorname{Span}_{\mathbb{R}}\{\,R_{\mathrm{off}}(X,Y)Z\;:\;X,Y,Z\in T_{p}M\,\}\ \subseteq\ \mathfrak{hol}_{p}(\nabla^{C}).$ (3.2)
Clearly $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$. If the action were reducible with respect to $V_{1}\oplus V_{2}$, all holonomy endomorphisms would be block-diagonal, forcing $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{0\}$. Hence nontrivial off-diagonal holonomy implies irreducibility.
4. Mixed Rank of a Degree-$3$ Class
Write the Künneth decomposition of $H^{3}(M)$ as:
$\big(H^{3}(M_{1})\otimes H^{0}(M_{2})\big)\ \oplus\ \big(H^{2}(M_{1})\otimes H^{1}(M_{2})\big)\ \oplus$ (4.1)
$\big(H^{1}(M_{1})\otimes H^{2}(M_{2})\big)\ \oplus\ \big(H^{0}(M_{1})\otimes H^{3}(M_{2})\big).$
Mixed components occur precisely in bidegrees $(2,1)$ and $(1,2)$.
Definition 4.1 (Mixed tensor rank).
Let $\omega]\in H^{3}(M;\mathbb{R})$. Denote by $\pi_{2,1}([\omega])\in H^{2}(M_{1})\otimes H^{1}(M_{2})$ and $\pi_{1,2}([\omega])\in H^{1}(M_{1})\otimes H^{2}(M_{2})$ its mixed Künneth projections. Define
$r_{2,1}=\min\Big\{r\,:\,\pi_{2,1}([\omega])=\sum_{i=1}^{r}\alpha_{i}\otimes\beta_{i},\ \alpha_{i}\in H^{2}(M_{1}),\ \beta_{i}\in H^{1}(M_{2})\Big\},$ (4.2)
and analogously $r_{1,2}$ for $\pi_{1,2}([\omega])$. The mixed rank is
$\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big):=r_{2,1}+r_{1,2}.$ (4.3)
For degree $3$, the tensor rank in each bidegree is well-defined and finite, and can be computed from any decomposition of the harmonic representative $\omega_{h}$ into wedge products of harmonic forms on the factors for the product metric $g$. Although $\omega_{h}$ depends on $g$, the ranks $r_{2,1},r_{1,2}$ are purely topological (indeed, the tensor rank of an element in $H^{2}(M_{1})\otimes H^{1}(M_{2})$ is defined as the minimal number of simple tensors in any decomposition, hence is invariant under change of metric within the product class).
5. Main Result and Proof
Lemma 5.1 (Generic rank of covariant derivatives).
Let $(M_{1},g_{1})$ be a closed oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms. Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$. Then the set
$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\text{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
is open and dense in $M_{1}$.
Proof.
Fix smooth local frames of $\Lambda^{2}T^{*}M_{1}$ and $T^{*}M_{1}$ over an open coordinate ball $B\subset M_{1}$. In these frames each $(\nabla^{g_{1}}\alpha_{i})(p_{1})$ becomes a smooth $\mathbb{R}$-valued matrix $A_{i}(p_{1})$ of size $\dim M_{1}\times\binom{\dim M_{1}}{2}$. Form the $(r_{2,1}-s)\times\bigl[\dim M_{1}\cdot\binom{\dim M_{1}}{2}\bigr$] matrix
$\mathbf{A}(p_{1})=\bigl[A_{s+1}(p_{1})^{\top}\!\!,\dots,A_{r_{2,1}}(p_{1})^{\top}\bigr]^{\top}.$
The condition $\text{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non-vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$. Each minor is a smooth function on $B$; hence its non-zero locus is open in $B$. Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open. Suppose, by contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non-empty open set $V\subset M_{1}$ on which
$\text{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
Thus, for every $p_{1}\in V$, there exists a non-zero vector $\lambda=(\lambda^{s+1},\dots,\lambda^{r_{2,1}})\in\mathbb{R}^{r_{2,1}-s}$ (depending on $p_{1}$) such that
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(\nabla^{g_{1}}\alpha_{i})(p_{1})=0.$
The smooth section
$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}\alpha_{i}\in\mathcal{H}^{2}(M_{1})$
then satisfies $\nabla^{g_{1}}\eta\equiv 0$ on $V$. By unique continuation for harmonic forms ([1]), a harmonic form parallel on an open set of any connected component of $M_{1}$ is globally parallel on that component (By Greene–Jacobowitz [4], we approximate $g$ in $C^{\infty}$ by real-analytic metrics $g_{k}$; for each such metric the Hodge-Laplace operator has analytic coefficients, so Aronszajn’s unique-continuation theorem applies to harmonic $2$-forms). Hence, $\eta$ is parallel on the connected component containing $V$, which implies $\eta\in\mathcal{P}_{2}$. This contradicts the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}$. Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
Theorem 5.2 (Topological lower bound for off-diagonal curvature).
Let $M=M_{1}\times M_{2}$ be compact and oriented, $g=g_{1}\oplus g_{2}$ any product metric, and $\nabla^{C}$ any cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]:=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
Then for every $p\in M$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\!\bigl(\mathcal{P}_{2}\otimes H^{1}(M_{2})\bigr),$
where $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ denotes the space of harmonic $2$-forms on $M_{1}$ that are parallel with respect to $g_{1}$.
The bound depends only on $\omega]_{\mathrm{mixed}}$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under $g_{1}$-deformations that preserve parallel $2$-forms, hence metric-invariant modulo parallel-form preserving deformations.
Proof.
Let $\omega_{h}$ be the $g$-harmonic representative of $\omega]$. Decompose $\omega_{h}$ according to Künneth and Hodge on the factors. For the $(2,1)$-part, write
$\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i},$ (5.1)
with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly, write
$\omega_{h}^{1,2}=\sum_{j=1}^{r_{1,2}}\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j},$ (5.2)
with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$. The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
Define the parallel subspace as
$\mathcal{P}_{2}:=\bigl\{\alpha\in\mathcal{H}^{2}(M_{1})\mid\nabla^{g_{1}}\alpha=0\bigr\}\subset H^{2}(M_{1};\mathbb{R})$ (5.3)
and set
$r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)-\dim\!\big(\mathcal{P}_{2}\otimes H^{1}(M_{2})\big).$ (5.4)
We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
Consider the curvature formula (2.5). The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources. First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$. Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off-diagonal operators. Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$. Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
To prove linear independence, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor-rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open-dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X,Y)|_{p}=\delta_{ik}$ and $\beta_{i}(Z)|_{p}=\delta_{il}$, obtaining $r$ linearly independent endomorphisms $R_{kl}^{\mathrm{off}}(p)\in\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. This produces at least $r$ linearly independent elements in $\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})$, since curvature endomorphisms generate the holonomy algebra by Ambrose–Singer [2].
For each simple summand $\alpha_{i}\otimes\beta_{i}]$ of the minimal tensor-rank decomposition, choose harmonic representatives $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$. By unique continuation for harmonic forms, their pointwise evaluations are linearly independent on an open-dense subset $\mathcal{U}\subset M$. For every $p=(p_{1},p_{2})\in\mathcal{U}$ the off-diagonal component of the linear part of the curvature,
$R^{\mathrm{off}}(X,Y)Z=\tfrac{1}{2}\big(\nabla^{LC}_{p}T\big)^{\mathrm{off}}(X,Y,Z),$
evaluated on vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$, defines a linear map
$\Psi_{p}:\mathrm{span}\{\alpha_{i}\otimes\beta_{i}\}\longrightarrow\mathfrak{so}(T_{p}M)$
whose injectivity follows from the pointwise independence of the evaluations $\{\alpha_{i}(p_{1})\}$ and $\{\beta_{i}(p_{2})\}$: any nontrivial relation among the images would imply a corresponding relation among the tensor products $\alpha_{i}(p_{1})\otimes\beta_{i}(p_{2})$, forcing all coefficients to vanish. Hence $\Psi_{p}$ produces $r$ linearly independent endomorphisms $R^{\mathrm{off}}(X_{i},Y_{i})Z_{i}$ in $\mathfrak{so}(T_{p}M)$.
The linear map
$\Psi_{p}:\mathrm{span}_{\mathbb{R}}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r}\to\mathfrak{so}(T_{p}M)$
defined by
$\Psi_{p}(\alpha_{i}\otimes\beta_{i})(Z):=\tfrac{1}{2}(\nabla^{LC}T)^{\mathrm{off}}(X,Y,Z)$
[…]
The map $\Psi_{p}$ is defined by the linear part of the curvature, which is proportional to $(\nabla^{LC}T)_{\text{off}}$. The condition that a linear combination of these contributions vanishes at a generic point $p=(p_{1},p_{2})$ is:
$\sum_{i}c_{i}\left[(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})\right]_{p}=0$
Since the forms $\alpha_{i}$ and $\beta_{i}$ are linearly independent and the covariant derivatives $\nabla^{g_{1}}$ and $\nabla^{g_{2}}$ act only on their respective factors, the vanishing of the mixed components of this sum implies that the terms must vanish separately.
Specifically, considering the component of type $(2,1)$ (forms with two indices on $M_{1}$ and one on $M_{2}$), the vanishing condition implies:
$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0$
Since the forms $\beta_{i}$ are linearly independent, this requires that the coefficients of $\beta_{i}$ must vanish:
$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})=0$
This condition holds on an open dense subset $\mathcal{U}\subset M$ (by Lemma 5.1). Since $\alpha_{i}$ are harmonic forms, the unique continuation principle implies that the linear combination $\sum c_{i}\alpha_{i}$ is a $\nabla^{g_{1}}$-parallel 2-form on $M_{1}$. By definition, the space of such forms is $\mathcal{P}_{2}$.
$\sum_{i}c_{i}\alpha_{i}\in\mathcal{P}_{2}$
This proves that the kernel of the map $\Psi_{p}$ is exactly the subspace generated by the simple tensors whose $\alpha$ component is $\nabla^{g_{1}}$-parallel.
Therefore, the kernel of the map $\Psi_{p}$ is precisely the intersection of the subspace spanned by the minimal decomposition with the space of parallel components:
$\ker\Psi_{p}=\left(\text{span}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r_{2,1}}\right)\cap\left(\mathcal{P}_{2}\otimes H^{1}(M_{2})\right)$
Since $\{\alpha_{i}\otimes\beta_{i}\}$ forms a basis for the mixed component $\pi_{2,1}([\omega])$, and the $\alpha_{i}$ are linearly independent, we can choose the basis such that the first $s=\dim(\mathcal{P}_{2})$ elements of $\{\alpha_{i}\}$ are a basis for $\mathcal{P}_{2}$. Given the linear independence of the $\beta_{i}$, the dimension of the kernel is precisely the number of simple tensors in the minimal decomposition whose $\alpha$-component is parallel. This dimension is $\dim(\ker\Psi_{p})=\dim\left(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}\otimes H^{1}(M_{2}))\right)$.
This establishes the dimension of the off-diagonal curvature generated by the linear term $(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$ on the open-dense subset $\mathcal{U}^{\prime}$ given by Lemma 5.1.
Consider the one-parameter family of connections
$\nabla^{(s)}=\nabla^{LC}+\frac{s}{2}T.$
Then the off-diagonal curvature satisfies $R_{\text{off}}(s)=s(\nabla^{LC}T)_{\text{off}}+s^{2}Q_{\text{off}}$. Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\text{off}}$ has dimension at least $r^{\sharp}$. The key step is to ensure that the rank of $R_{\text{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\text{off}}$.
We consider the one-parameter family of connections $\nabla^{(s)}=\nabla^{LC}+s\frac{1}{2}T$. The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$. On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$.
To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in the Remark 5.6).
By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), there exists a sequence of real-analytic metrics $g_{k}\to g$ in the $C^{\infty}$-topology such that $\dim\mathcal{P}_{2}(g_{k})=\dim\mathcal{P}_{2}(g)$ for every $k$; this is possible because the stratum $\{g^{\prime}\mid\dim\mathcal{P}_{2}(g^{\prime})=\dim\mathcal{P}_{2}(g)\}$ is open in the $C^{\infty}$-topology and contains real-analytic metrics densely (Nash–Moser, [9]). Consequently the curvature $R_{\mathrm{off},k}(s)$ associated with $g_{k}$ is real-analytic in $(s,p)$ and the integer $r^{\sharp}$ is the same for every $g_{k}$.
For each fixed $k$, the matrix $R_{\mathrm{off},k}(s)$ is analytic. Since $\mathrm{rank}(A_{1,k})=r^{\sharp}$ on $U$, there exists an $r^{\sharp}\times r^{\sharp}$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ such that $M_{k}(s)=s^{r^{\sharp}}D_{k}(s)$, where $D_{k}(0)\neq 0$. By the Identity Principle for Analytic Functions, $D_{k}(s)$ is not identically zero, and its set of zeros is discrete. Thus, $\mathrm{rank}\,R_{\mathrm{off},k}(p)\geq r^{\sharp}$ holds for all $p\in U$ and for all $k$.
The rank function is lower-semicontinuous with respect to $C^{0}$ convergence ([10]). Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, the inequality is preserved:
$\mathrm{rank}(R_{\mathrm{off}}(p))\geq\liminf_{k\to\infty}\mathrm{rank}(R_{\mathrm{off},k}(p))\geq r^{\sharp}\quad\text{on a dense subset of }M.$
The extension of the inequality to the entire manifold $M$ follows from the fact that the set $\{p\in M\mid\mathrm{rank}\,R_{\mathrm{off}}(p)\geq r^{\sharp}\}$ is open (by lower-semicontinuity) and contains a dense subset, thus coinciding with $M$.
To extend the inequality
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$
from the dense open subset $\mathcal{U}$ to all of $M$, we use the lower-semicontinuity of the rank of a continuous family of linear maps.
The extension of the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ to the entire manifold $M$ follows from the lower-semicontinuity of the rank, as established in [10]. Indeed, by Ambrose–Singer the holonomy algebra is generated by the curvature operators; since the rank of a continuous family of Lie-algebra generators is lower-semicontinuous (Warner [10]), the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ extends from the dense open set to all of $M$.
Finally, the lower bound depends only on $\omega]$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under any deformation of $g_{1}$ that preserves the space of parallel $2$-forms, hence under any product-metric deformation that leaves $\mathcal{P}_{2}$ unchanged. ∎
Remark 5.3 (Necessity of the Non-Parallel Torsion Condition).
The proof of Theorem 5.1 relies fundamentally on the non-vanishing of the linear term $R^{C}_{\mathrm{off}}\simeq\frac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$, which is established via the $s\to 0$ limit of the one-parameter family of connections $\nabla^{(s)}$. This dependence on the linear term necessitates the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof to be valid.
To illustrate the necessity of this condition, we consider a case where $\nabla^{\mathrm{LC}}T=0$. If $\nabla^{\mathrm{LC}}T=0$, the curvature formula simplifies to $R^{C}=R^{\mathrm{LC}}+\frac{1}{4}Q_{T}$, where $Q_{T}$ is the term quadratic in the torsion $T$. Since $R^{\mathrm{LC}}$ is block-diagonal for a product metric, the off-diagonal curvature is solely determined by $Q_{T}^{\mathrm{off}}$. The proof of Theorem 5.1 provides no guarantee that the quadratic term $Q_{T}^{\mathrm{off}}$ alone is sufficient to generate $r^{\sharp}$ independent directions.
Consider the flat torus $M=T^{3}=T^{2}\times S^{1}$ endowed with the flat product metric $g$. Let $T$ be the constant volume form $T=dx\wedge dy\wedge dz$. This form is harmonic, non-trivial, and its mixed rank is $\text{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1$. Since $T$ is constant, we have $\nabla^{\mathrm{LC}}T=0$.
For this specific example, the curvature $R^{C}$ vanishes identically, leading to $\dim(\text{hol}_{p}^{\mathrm{off}}(\nabla^{C}))=0$. The lower bound is calculated as $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2}))=1-1=0$.
The resulting inequality $0\geq 0$ is satisfied, but the bound is trivial. This example demonstrates that when $\nabla^{\mathrm{LC}}T=0$, the topological lower bound $r^{\sharp}$ can be reduced to zero, even for a non-trivial cohomology class ($\omega]\neq 0$). The failure of the bound to provide a non-trivial constraint confirms that the contribution of the quadratic term $Q_{T}^{\mathrm{off}}$ alone is insufficient to guarantee the non-triviality of the topological lower bound, necessitating the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof of Theorem 5.1 to hold and to yield a meaningful geometric result.  [^1] The bound is recovered if we consider a non-parallel torsion $T=\omega_{h}+\varepsilon d(\sin(x)dy\wedge dz)$, which maintains the same cohomology class $\omega_{h}]$ but ensures $\nabla^{\mathrm{LC}}T\neq 0$.
Remark 5.4 (Geometric Significance of Invariance).
The invariance of the lower bound under $g_{1}$-deformations that preserve $\mathcal{P}_{2}$ is non-trivial and geometrically significant on manifolds $M_{1}$ with special holonomy (e.g., Calabi-Yau or flat tori). In these cases, the space of parallel forms $\mathcal{P}_{2}$ is non-zero, and the existence of non-homothetic deformations preserving $\mathcal{P}_{2}$ is guaranteed by known results in differential geometry (e.g., deformation theory of special holonomy metrics). This highlights the topological nature of the bound, which is stable under a non-trivial class of metric variations.
Remark 5.5 (Forced irreducibility and quantified off-diagonal holonomy).
Since $\nabla^{C}$ is metric, any non-trivial off-diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$. Theorem 5.2 quantifies the “forced irreducibility” principle in terms of the mixed tensor rank, modulo the parallel subspace $\mathcal{P}_{2}\otimes H^{1}(M_{2})$. Specifically, the dimension of the off-diagonal holonomy subspace is bounded below by
$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$. Irreducibility is forced if and only if $r^{\sharp}>0$, i.e. when the mixed tensor rank exceeds the dimension of the parallel subspace.
Remark 5.6 (Analyticity as a proof device).
The use of real-analytic metrics in the proof of Theorem 5.2 is purely instrumental: analytic approximation (Greene–Jacobowitz) and lower-semicontinuity of matrix rank extend the bound to every $C^{\infty}$ product metric without additional assumptions.
6. Examples and Sharpness
The case $M=S^{2}\times\Sigma_{g}$
Let $\Sigma_{g}$ be a closed oriented surface of genus $g\geq 1$. By Künneth,
$H^{3}(S^{2}\times\Sigma_{g};\mathbb{R})\cong H^{2}(S^{2})\otimes H^{1}(\Sigma_{g}),$ (6.1)
since $H^{1}(S^{2})=0$ and $H^{2}(\Sigma_{g})\cong\mathbb{R}$. The mixed component is therefore isomorphic to $\mathbb{R}^{2g}$. Choosing a harmonic basis $\{\alpha_{j}\}_{j=1}^{2g}$ for $H^{1}(\Sigma_{g})$ and the volume form $\mathrm{vol}_{S^{2}}$ on $S^{2}$, any mixed class has a harmonic representative of the form
$\omega_{h}=\sum_{j=1}^{2g}c_{j}\,\mathrm{vol}_{S^{2}}\wedge\alpha_{j}.$ (6.2)
Thus $\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)=2g$. For each $j$, select $X,Y\in V_{1}$ so that $\mathrm{vol}_{S^{2}}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\alpha_{j}(Z)\neq 0$; the corresponding curvature endomorphisms produce $2g$ independent off-diagonal directions. Hence
$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2g-\dim(\mathcal{P}_{2}\otimes H^{1}(\Sigma_{g})),$ (6.3)
where $\mathcal{P}_{2}\subset H^{2}(S^{2};\mathbb{R})$ is the subspace of harmonic parallel $2$-forms on $S^{2}$ (here $\mathcal{P}_{2}=\{0\}$, so the bound is $2g$).
The case $M=S^{2}\times T^{2}$
Here $H^{3}(M)\cong H^{2}(S^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2}$. Writing $\omega_{h}=c_{1}\,\mathrm{vol}_{S^{2}}\wedge dx+c_{2}\,\mathrm{vol}_{S^{2}}\wedge dy$ with $dx,dy$ the harmonic $1$-forms on the torus, one obtains two independent mixed components, hence
$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2-\dim(\mathcal{P}_{2}\otimes H^{1}(T^{2})),$ (6.4)
where $\mathcal{P}_{2}=\{0\}$ on $S^{2}$, so the bound is $2$.
7. Concluding Remarks and Perspectives
We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with totally skew torsion calibrated by a mixed cohomology class, the off-diagonal holonomy dimension admits a quantified lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ is the space of harmonic parallel $2$-forms on $M_{1}$. The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds. This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace. The examples above demonstrate sharpness in natural settings where the torsion equals the harmonic representative and the parallel subspace has minimal dimension. It is natural to conjecture that equality holds whenever the torsion is purely harmonic, whereas strict inequality may occur when the coexact or exact parts of $T$ generate additional independent off-diagonal curvature directions or when the parallel subspace has maximal dimension. A systematic classification of such cases is an interesting direction for future work. A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].
Compactness and Hodge theory. Compactness ensures the existence and uniqueness of harmonic representatives [10], and the vanishing of boundary terms in integrations by parts. Extending the results to noncompact manifolds would require an $L^{2}$ Hodge framework or alternative analytic hypotheses.

### v9

@@
 Abstract.
-We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of cohomologically calibrated metric connections with totally skew torsion on product manifolds.
-We show that the dimension of the off-diagonal curvature subspace of a connection $\nabla^{C}$ is bounded below by the sum of tensor ranks of the mixed Künneth components of its calibration class, minus the dimension of the space of harmonic parallel forms on the first factor.
-The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}\in H^{3}(M;\mathbb{R})$ and on the space of $g_{1}$-parallel harmonic $2$-forms on $M_{1}$; it is therefore independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves parallel $2$-forms, hence it is metric-invariant modulo parallel-form preserving deformations.
-This provides a computational criterion for geometric complexity and quantifies the interaction between topology and curvature, yielding a quantified version of “forced irreducibility” via the dimension of $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
+We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of metric connections with totally skew torsion on product manifolds.
+For any such connection whose non‑parallel torsion 33‑form represents a mixed de Rham class via its harmonic projection, we prove that the dimension of the off‑diagonal holonomy subspace satisfies the lower bound
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
+where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the linear obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$.
+These kernels are defined intrinsically as the intersections
+$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
+where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (respectively $\mathcal{V}_{1,2}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$) denotes the linear span of the harmonic representatives of the $(2,1)$- (respectively $(1,2)$-) Künneth component of $\omega]$.
+The subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are independent of the choice of a minimal harmonic decomposition.
+Consequently, the integer $\dim\mathcal{K}$ depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics; hence the bound is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
+This yields a computable, metric‑invariant obstruction to reducible holonomy and quantifies the topological “mixing strength” that forces irreducibility across the product splitting.
 Keywords: Connections with Torsion, Product Manifolds, Holonomy Algebra, De Rham Cohomology, Hodge Theory, Künneth Decomposition, Curvature Tensor
@@
 Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$.
-We introduce a topological quantity, the mixed tensor rank of the cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that it furnishes, for any metric connection $\nabla^{\mathcal{C}}$, a quantified lower bound for the dimension of the subspace of the holonomy algebra generated by off-diagonal curvature operators with respect to the splitting $V_{1}\oplus V_{2}$.
-The key point is that, although the definition of “off-diagonal” uses $g$, the lower bound is independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves the space of parallel 2-forms, and is explicitly corrected by the dimension of $g_{1}$-parallel harmonic 2-forms on the first factor.
-Each independent mixed Künneth component in $\omega]$ guarantees a linearly independent off-diagonal curvature direction, modulo the parallel subspace.
+We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
+More precisely, we prove
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
+where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the natural linear obstruction maps. To define these kernels, let
+$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
+where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp.
+$\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp.
+$(1,2)$-) Künneth component of $\omega]$.
+Because any two minimal harmonic decompositions span the same subspaces $U\subset\mathcal{H}^{2}(M_{1})$ and $V\subset\mathcal{H}^{1}(M_{2})$, the spaces $\mathcal{V}_{2,1}=U\otimes V$ and $\mathcal{V}_{1,2}$ are well‑defined, i.e., independent of the chosen decomposition.
+Then
+$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
+The integer $\dim\mathcal{K}$ measures precisely the part of the mixed class that is absorbed by the parallel‑form strata and therefore does not contribute to off‑diagonal curvature.
+The quantity $r^{\sharp}$ is metric‑invariant under deformations that preserve the parallel‑form strata: it depends only on $\omega]_{\mathrm{mixed}}$ and on the subspaces $\mathcal{P}_{k}(M_{i})$.
+Consequently, each independent mixed Künneth tensor guarantees a linearly independent off‑diagonal holonomy direction, modulo the parallel forms captured by $\mathcal{K}$.
+This yields a computable, purely topological criterion for forced irreducibility of the holonomy representation across the product splitting: irreducibility is forced whenever $r^{\sharp}>0$, i.e.
+whenever $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}$.
 2. Cohomologically Calibrated Metric Connections with Skew Torsion
@@
 Definition 2.1 (Metric cohomological calibration).
-A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the cohomology class $\omega]$ is non trivial and defined via the harmonic projection of the torsion three-form, ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, so $\omega]:=[\omega_{h}]$), is called cohomologically calibrated.
+A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the torsion 3-form represents a non trivial mixed cohomology class $\omega]$ via its harmonic projection ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, i.e.
+$\omega]:=[\omega_{h}]$), is called cohomologically calibrated.
 The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
@@
 Definition 3.2 (Off-diagonal holonomy subspace).
-The off-diagonal holonomy subspace at $p\in M$ is
-$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\operatorname{Span}_{\mathbb{R}}\{\,R_{\mathrm{off}}(X,Y)Z\;:\;X,Y,Z\in T_{p}M\,\}\ \subseteq\ \mathfrak{hol}_{p}(\nabla^{C}).$ (3.2)
-Clearly $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
+Let $(M,g)$ be a Riemannian manifold and $\nabla^{C}$ a metric connection on $TM$ with totally skew–symmetric torsion.
+For a point $p\in M$, we define the off–diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ as the Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ generated by all off–diagonal components of the curvature tensor and its covariant derivatives at $p$, namely
+$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C}):=\mathrm{Lie}\,\Big\langle\big(\nabla^{k}R^{C}\big)_{p}(X_{1},\ldots,X_{k};Y,Z)^{\mathrm{off}}\;\Big|\;k\geq 0,\;X_{i},Y,Z\in T_{p}M\Big\rangle,$
+where $(\cdot)^{\mathrm{off}}$ denotes the off–diagonal projection with respect to the product splitting $T_{p}M=V_{1}\oplus V_{2}$.
+By the Ambrose–Singer theorem [2], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor.
+In particular, at each point $p\in M$ one has, clearly, $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
 Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$.
@@
 Lemma 5.1 (Generic rank of covariant derivatives).
-Let $(M_{1},g_{1})$ be a closed oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms.
+Let $(M_{1},g_{1})$ be a compact and oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms.
 Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$.
 Then the set
-$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\text{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
+$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\operatorname{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
 is open and dense in $M_{1}$.
@@
 $\mathbf{A}(p_{1})=\bigl[A_{s+1}(p_{1})^{\top}\!,\dots,A_{r_{2,1}}(p_{1})^{\top}\bigr]^{\top}.$
-The condition $\text{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non-vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$.
-Each minor is a smooth function on $B$; hence its non-zero locus is open in $B$.
+The condition $\operatorname{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non‑vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$.
+Each minor is a smooth function on $B$; hence its non‑zero locus is open in $B$.
 Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open.
-Suppose, by contradiction, that $\mathcal{U}^{\prime}$ is not dense.
-Then there exists a non-empty open set $V\subset M_{1}$ on which
-$\text{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
-Thus, for every $p_{1}\in V$, there exists a non-zero vector $\lambda=(\lambda^{s+1},\dots,\lambda^{r_{2,1}})\in\mathbb{R}^{r_{2,1}-s}$ (depending on $p_{1}$) such that
-$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(\nabla^{g_{1}}\alpha_{i})(p_{1})=0.$
-The smooth section
-$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}\alpha_{i}\in\mathcal{H}^{2}(M_{1})$
-then satisfies $\nabla^{g_{1}}\eta\equiv 0$ on $V$.
-By unique continuation for harmonic forms ([1]), a harmonic form parallel on an open set of any connected component of $M_{1}$ is globally parallel on that component (By Greene–Jacobowitz [4], we approximate $g$ in $C^{\infty}$ by real-analytic metrics $g_{k}$; for each such metric the Hodge-Laplace operator has analytic coefficients, so Aronszajn’s unique-continuation theorem applies to harmonic $2$-forms).
-Hence, $\eta$ is parallel on the connected component containing $V$, which implies $\eta\in\mathcal{P}_{2}$.
-This contradicts the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}$.
+Now suppose, for a contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non‑empty open set $V\subset M_{1}$ on which
+$\operatorname{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
+Since the rank is lower‑semicontinuous, after possibly shrinking $V$ we may assume that the rank is constant, say $\rho<r_{2,1}-s$, on $V$.
+Consequently, there exists a smooth map $\lambda:V\to\mathbb{R}^{r_{2,1}-s}\setminus\{0\}$ such that
+$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{1})(\nabla^{g_{1}}\alpha_{i})(p_{1})=0\qquad\forall p_{1}\in V.$ (5.1)
+Now we employ analytic approximation.
+By the theorem of Greene–Jacobowitz [4], we can approximate the metric $g_{1}$ in the $C^{\infty}$ topology by a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}$.
+For each $k$, denote by $\alpha_{i}^{(k)}$ the $g_{1}^{(k)}$-harmonic representative of the cohomology class $\alpha_{i}]$.
+The forms $\alpha_{i}^{(k)}$ converge to $\alpha_{i}$ in $C^{\infty}$ as $k\to\infty$.
+Because $\lambda$ is smooth and non‑zero on $V$, we may choose a point $p_{0}\in V$ and a subsequence (still denoted by $k$) such that $\lambda(p_{0})\neq 0$ and
+$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})(\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)})(p_{0})=0\qquad\text{for all }k.$ (5.2)
+Indeed, (5.1) passes to the limit after extracting a suitable subsequence, since all objects depend smoothly on the metric.
+For each fixed $k$, the metric $g_{1}^{(k)}$ is real‑analytic. Define
+$\eta^{(k)}:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}^{(k)}\in\mathcal{H}^{2}(M_{1},g_{1}^{(k)}).$
+Equation (5.2) states that $\nabla^{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$.
+Because $g_{1}^{(k)}$ is analytic, the coefficients of the Hodge–Laplacian are analytic, and Aronszajn’s unique-continuation theorem [1] applies.
+Since $\eta^{(k)}$ is harmonic and $g_{1}^{(k)}$ is real-analytic, $\eta^{(k)}$ is also real-analytic.
+The condition $\nabla_{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$ implies that $\eta^{(k)}$ is parallel in a neighborhood of $p_{0}$.
+By the unique-continuation principle for parallel forms on real-analytic manifolds (a consequence of Aronszajn’s theorem applied to the analytic coefficients), $\eta^{(k)}$ must be parallel on the entire connected component.
+Passing to the limit $k\to\infty$, we obtain that
+$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}$
+is parallel with respect to the original metric $g_{1}$ on the component containing $p_{0}$.
+Thus $\eta\in\mathcal{P}_{2}(M_{1})$, contradicting the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}(M_{1})$.
 Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
+The same reasoning applies verbatim to $\mathcal{H}^{k}(M_{i})$ for any $k$ and for both factors $M_{1}$ and $M_{2}$. ∎
 Theorem 5.2 (Topological Lower Bound for Off-Diagonal Curvature).
-Let $M=M_{1}\times M_{2}$ be compact and oriented, $g=g_{1}\oplus g_{2}$ any product metric, and $\nabla^{C}$ any cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]:=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
-Then for every $p\in M$,
-$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\!\bigl(\mathcal{P}_{2}\otimes H^{1}(M_{2})\bigr),$
-where $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ denotes the space of harmonic $2$-forms on $M_{1}$ that are parallel with respect to $g_{1}$.
-The bound depends only on $\omega]_{\mathrm{mixed}}$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under $g_{1}$-deformations that preserve parallel $2$-forms, hence metric-invariant modulo parallel-form preserving deformations.
+Let $M=M_{1}\times M_{2}$ be a compact and oriented product manifold with a product metric $g=g_{1}\oplus g_{2}$, and let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
+Let
+$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
+where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp.
+$\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp.
+$(1,2)$-) Künneth component of $\omega]$.
+Define
+$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
+and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$. Then, for every $p\in M$,
+$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
+The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
 Proof.
@@
 $\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i},$ (5.3)
-with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly, write
+with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly,
 $\omega_{h}^{1,2}=\sum_{j=1}^{r_{1,2}}\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j},$ (5.4)
 with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$.
-The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
-Define the parallel subspace as
-$\mathcal{P}_{2}:=\bigl\{\alpha\in\mathcal{H}^{2}(M_{1})\mid\nabla^{g_{1}}\alpha=0\bigr\}\subset H^{2}(M_{1};\mathbb{R})$ (5.3)
-and set
-$r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)-\dim\!\big(\mathcal{P}_{2}\otimes H^{1}(M_{2})\big).$ (5.4)
-We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
+The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
+Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
 Consider the curvature formula (2.5).
@@
 First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$.
-Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off-diagonal operators.
+Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off‑diagonal operators.
 Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$.
 Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
-To prove linear independence, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$.
-By minimality of the tensor-rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open-dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent.
-For any such $p$ we pick vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X,Y)|_{p}=\delta_{ik}$ and $\beta_{i}(Z)|_{p}=\delta_{il}$, obtaining $r$ linearly independent endomorphisms $R_{kl}^{\mathrm{off}}(p)\in\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
-This produces at least $r$ linearly independent elements in $\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})$, since curvature endomorphisms generate the holonomy algebra by Ambrose–Singer [2].
-For each simple summand $\alpha_{i}\otimes\beta_{i}]$ of the minimal tensor-rank decomposition, choose harmonic representatives $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$.
-By unique continuation for harmonic forms, their pointwise evaluations are linearly independent on an open-dense subset $\mathcal{U}\subset M$.
-For every $p=(p_{1},p_{2})\in\mathcal{U}$ the off-diagonal component of the linear part of the curvature,
-$R^{\mathrm{off}}(X,Y)Z=\tfrac{1}{2}\big(\nabla^{LC}_{p}T\big)^{\mathrm{off}}(X,Y,Z),$
-evaluated on vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$, defines a linear map
-$\Psi_{p}:\mathrm{span}\{\alpha_{i}\otimes\beta_{i}\}\longrightarrow\mathfrak{so}(T_{p}M)$
-whose injectivity follows from the pointwise independence of the evaluations $\{\alpha_{i}(p_{1})\}$ and $\{\beta_{i}(p_{2})\}$: any nontrivial relation among the images would imply a corresponding relation among the tensor products $\alpha_{i}(p_{1})\otimes\beta_{i}(p_{2})$, forcing all coefficients to vanish.
-Hence $\Psi_{p}$ produces $r$ linearly independent endomorphisms $R^{\mathrm{off}}(X_{i},Y_{i})Z_{i}$ in $\mathfrak{so}(T_{p}M)$.
-The linear map
-$\Psi_{p}:\mathrm{span}_{\mathbb{R}}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r}\to\mathfrak{so}(T_{p}M)$
-defined by
-$\Psi_{p}(\alpha_{i}\otimes\beta_{i})(Z):=\tfrac{1}{2}(\nabla^{LC}T)^{\mathrm{off}}(X,Y,Z)$
-[…]
-The map $\Psi_{p}$ is defined by the linear part of the curvature, which is proportional to $(\nabla^{LC}T)_{\text{off}}$.
-The condition that a linear combination of these contributions vanishes at a generic point $p=(p_{1},p_{2})$ is:
-$\sum_{i}c_{i}\left[(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})\right]_{p}=0$
-Since the forms $\alpha_{i}$ and $\beta_{i}$ are linearly independent and the covariant derivatives $\nabla^{g_{1}}$ and $\nabla^{g_{2}}$ act only on their respective factors, the vanishing of the mixed components of this sum implies that the terms must vanish separately.
-Specifically, considering the component of type $(2,1)$ (forms with two indices on $M_{1}$ and one on $M_{2}$), the vanishing condition implies:
-$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0$
-Since the forms $\beta_{i}$ are linearly independent, this requires that the coefficients of $\beta_{i}$ must vanish:
-$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})=0$
-This condition holds on an open dense subset $\mathcal{U}\subset M$ (by Lemma 5.1).
-Since $\alpha_{i}$ are harmonic forms, the unique continuation principle implies that the linear combination $\sum c_{i}\alpha_{i}$ is a $\nabla^{g_{1}}$-parallel 2-form on $M_{1}$.
-By definition, the space of such forms is $\mathcal{P}_{2}$.
-$\sum_{i}c_{i}\alpha_{i}\in\mathcal{P}_{2}$
-This proves that the kernel of the map $\Psi_{p}$ is exactly the subspace generated by the simple tensors whose $\alpha$ component is $\nabla^{g_{1}}$-parallel.
-Therefore, the kernel of the map $\Psi_{p}$ is precisely the intersection of the subspace spanned by the minimal decomposition with the space of parallel components:
-$\ker\Psi_{p}=\left(\text{span}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r_{2,1}}\right)\cap\left(\mathcal{P}_{2}\otimes H^{1}(M_{2})\right)$
-Since $\{\alpha_{i}\otimes\beta_{i}\}$ forms a basis for the mixed component $\pi_{2,1}([\omega])$, and the $\alpha_{i}$ are linearly independent, we can choose the basis such that the first $s=\dim(\mathcal{P}_{2})$ elements of $\{\alpha_{i}\}$ are a basis for $\mathcal{P}_{2}$.
-Given the linear independence of the $\beta_{i}$, the dimension of the kernel is precisely the number of simple tensors in the minimal decomposition whose $\alpha$-component is parallel.
-This dimension is $\dim(\ker\Psi_{p})=\dim\left(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}\otimes H^{1}(M_{2}))\right)$.
-This establishes the dimension of the off-diagonal curvature generated by the linear term $(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$ on the open-dense subset $\mathcal{U}^{\prime}$ given by Lemma 5.1.
+To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$.
+By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open‑dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent.
+For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
+Define the linear map
+$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}).$ (5.5)
+The map is well‑defined and linear; its image consists precisely of the off‑diagonal curvature operators generated by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the (2,1)(2,1)‑component of $\omega_{h}$.
+If $\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$, then the corresponding off‑diagonal curvature contribution vanishes:
+$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
+Using the product rule $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})$ and the fact that $R^{LC}_{\mathrm{off}}=0$, this condition is equivalent to
+$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0.$ (5.6)
+Separating the bidegrees in (5.6) yields the two independent conditions
+$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.7)
+Because the forms $\beta_{i}$ (resp. $\alpha_{i}$) are linearly independent on an open‑dense set, we obtain
+$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}=0$ (5.8)
+on that set. By the unique‑continuation principle for harmonic forms (Lemma 5.1), these equalities extend globally, hence
+$\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.9)
+Conversely, if (5.9) holds, then $\nabla^{LC}\bigl(\sum_{i}c_{i}\alpha_{i}\wedge\beta_{i}\bigr)=0$, which forces $\Psi_{p}\bigl(\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\bigr)=0$.
+Therefore $\xi=\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}$ belongs to $\ker\Psi_{p}$ if and only if (5.9) is satisfied.
+Now we identify this kernel with the intersection $\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)$.
+If $\xi\in\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}\otimes\mathcal{P}_{1}\bigr)$, then $\xi$ can be written as a finite sum $\sum_{k}A_{k}\otimes B_{k}$ with $A_{k}\in\mathcal{P}_{2}$, $B_{k}\in\mathcal{P}_{1}$.
+For each term, $\nabla^{LC}(A_{k}\wedge B_{k})=0$ because $A_{k}$ and $B_{k}$ are parallel, hence $\Psi_{p}(A_{k}\otimes B_{k})=0$.
+By linearity, $\Psi_{p}(\xi)=0$, so $\xi\in\ker\Psi_{p}$.
+For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Equation (5.9) gives
+$A:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad B:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
+Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$.
+Set $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and $V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2})$.
+Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{\dim U}\}$ whose classes are a basis of the quotient $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$).
+By Lemma 5.1, at a generic point $p_{1}\in M_{1}$ the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{\dim U}$ are linearly independent.
+Since $A$ is parallel, $\nabla^{g_{1}}A\equiv 0$; evaluating at such a $p_{1}$ yields
+$\sum_{i>s}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0,$
+and the linear independence forces $c_{i}=0$ for all $i>s$.
+An identical construction applied to $M_{2}$ (choosing a basis of $V_{\parallel}$ and completing it to a basis of $V$ such that the remaining vectors form a basis of the quotient $V/V_{\parallel}$) implies $c_{i}=0$ for all $i>t$.
+Hence, the only possibly non-zero coefficients are those with $i\leq\min(s,t)$.
+For such indices, by construction $\alpha_{i}\in\mathcal{P}_{2}(M_{1})$ and $\beta_{i}\in\mathcal{P}_{1}(M_{2})$.
+Therefore,
+$\xi=\sum_{i=1}^{\min(s,t)}c_{i}\alpha_{i}\otimes\beta_{i}\in\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}).$
+Since $\xi\in\mathcal{V}_{2,1}$ by hypothesis, we have shown
+$\ker\Psi_{p}\subseteq\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
+The opposite inclusion is immediate: any element of $\mathcal{V}_{2,1}$ that belongs to $\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})$ is a finite sum of simple tensors of parallel forms; consequently its covariant derivative vanishes, and hence it lies in $\ker\Psi_{p}$.
+Thus we have proved
+$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
+An identical argument applied to the $(1,2)$-component yields
+$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
+Now consider the induced map on the quotient
+$\overline{\Psi}_{p}:\mathcal{V}_{2,1}/\ker\Psi_{p}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\overline{\Psi}_{p}([\xi]):=\Psi_{p}(\xi).$
+The map $\overline{\Psi}_{p}$ is well‑defined and linear.
+It is injective because $\overline{\Psi}_{p}([\xi])=0$ implies $\Psi_{p}(\xi)=0$, hence $\xi\in\ker\Psi_{p}$ and $\xi]=0$.
+Consequently, $\overline{\Psi}_{p}$ sends a basis of the quotient to a set of linearly independent off‑diagonal curvature operators, and we obtain
+$\dim(\operatorname{im}\Psi_{p})=\dim\bigl(\mathcal{V}_{2,1}/\ker\Psi_{p}\bigr)=r_{2,1}-\dim\ker\Psi_{p}.$ (5.10)
+An entirely analogous argument for the (1,2)(1,2)‑component gives
+$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)$ (5.11)
+and
+$\dim(\operatorname{im}\widetilde{\Psi}_{p})=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$ (5.12)
+Hence, on the open‑dense set $\mathcal{U}$ where the harmonic evaluations are independent, the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
+$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
+linearly independent off‑diagonal curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra, whence
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
 Consider the one-parameter family of connections
 $\nabla^{(s)}=\nabla^{LC}+\frac{s}{2}T.$
-Then the off-diagonal curvature satisfies $R_{\text{off}}(s)=s(\nabla^{LC}T)_{\text{off}}+s^{2}Q_{\text{off}}$.
-Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\text{off}}$ has dimension at least $r^{\sharp}$.
-The key step is to ensure that the rank of $R_{\text{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\text{off}}$.
-We consider the one-parameter family of connections $\nabla^{(s)}=\nabla^{LC}+s\frac{1}{2}T$.
+Then the off-diagonal curvature satisfies $R_{\mathrm{off}}(s)=s(\nabla^{LC}T)_{\mathrm{off}}+s^{2}Q_{\mathrm{off}}$.
+Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\mathrm{off}}$ has dimension at least $r^{\sharp}$.
+The key step is to ensure that the rank of $R_{\mathrm{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\mathrm{off}}$.
 The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$.
 On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$.
-To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in the Remark 5.6).
-By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), there exists a sequence of real-analytic metrics $g_{k}\to g$ in the $C^{\infty}$-topology such that $\dim\mathcal{P}_{2}(g_{k})=\dim\mathcal{P}_{2}(g)$ for every $k$; this is possible because the stratum $\{g^{\prime}\mid\dim\mathcal{P}_{2}(g^{\prime})=\dim\mathcal{P}_{2}(g)\}$ is open in the $C^{\infty}$-topology and contains real-analytic metrics densely (Nash–Moser, [9]).
-Consequently the curvature $R_{\mathrm{off},k}(s)$ associated with $g_{k}$ is real-analytic in $(s,p)$ and the integer $r^{\sharp}$ is the same for every $g_{k}$.
-For each fixed $k$, the matrix $R_{\mathrm{off},k}(s)$ is analytic.
-Since $\mathrm{rank}(A_{1,k})=r^{\sharp}$ on $U$, there exists an $r^{\sharp}\times r^{\sharp}$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ such that $M_{k}(s)=s^{r^{\sharp}}D_{k}(s)$, where $D_{k}(0)\neq 0$.
-By the Identity Principle for Analytic Functions, $D_{k}(s)$ is not identically zero, and its set of zeros is discrete.
-Thus, $\mathrm{rank}\,R_{\mathrm{off},k}(p)\geq r^{\sharp}$ holds for all $p\in U$ and for all $k$.
-The rank function is lower-semicontinuous with respect to $C^{0}$ convergence ([10]). Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, the inequality is preserved:
-$\mathrm{rank}(R_{\mathrm{off}}(p))\geq\liminf_{k\to\infty}\mathrm{rank}(R_{\mathrm{off},k}(p))\geq r^{\sharp}\quad\text{on a dense subset of }M.$
-The extension of the inequality to the entire manifold $M$ follows from the fact that the set $\{p\in M\mid\mathrm{rank}\,R_{\mathrm{off}}(p)\geq r^{\sharp}\}$ is open (by lower-semicontinuity) and contains a dense subset, thus coinciding with $M$.
-To extend the inequality
-$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$
-from the dense open subset $\mathcal{U}$ to all of $M$, we use the lower-semicontinuity of the rank of a continuous family of linear maps.
-The extension of the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ to the entire manifold $M$ follows from the lower-semicontinuity of the rank, as established in [10].
-Indeed, by Ambrose–Singer the holonomy algebra is generated by the curvature operators; since the rank of a continuous family of Lie-algebra generators is lower-semicontinuous (Warner [10]), the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ extends from the dense open set to all of $M$.
-Finally, the lower bound depends only on $\omega]$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under any deformation of $g_{1}$ that preserves the space of parallel $2$-forms, hence under any product-metric deformation that leaves $\mathcal{P}_{2}$ unchanged.
+The rank $r^{\sharp}$ is established by the harmonic component $\omega_{h}$ of $T$.
+The non-harmonic part $T-\omega_{h}=d\alpha+\delta\beta$ contributes to $A_{1}$, but since the bound is a lower bound, this contribution cannot reduce the rank below $r^{\sharp}$ on the dense set $U$ and is therefore absorbed by the overall estimate $\mathrm{rank}(A_{1})\geq r^{\sharp}$.
+More formally, a systematic cancellation of the off-diagonal contributions by the non-harmonic part $\eta=d\alpha+\delta\beta$ on an open set would require $\nabla\eta=-\nabla\omega_{h}$ locally.
+By the Unique Continuation Principle, this identity would extend globally, implying $\omega_{h}]=0$, which contradicts our hypothesis of a non-trivial mixed class.
+This ensures that the topological information carried by $\omega_{h}$ is rigidly encoded in the curvature.
+To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in Remark 5.6).
+By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), we can select a sequence of real-analytic product metrics $g_{k}=g_{1,k}\oplus g_{2,k}$ converging to $g$ in the $C^{\infty}$ topology and satisfying
+$\dim\mathcal{P}_{k}(M_{i},g_{k})\leq\dim\mathcal{P}_{k}(M_{i},g)\qquad(k=1,2).$
+This selection is possible because the dimension function $g^{\prime}\mapsto\dim\mathcal{P}_{k}(M_{i},g^{\prime})$ is upper semicontinuous; consequently, the set $\{g^{\prime}\mid\dim\mathcal{P}_{k}(M_{i},g^{\prime})\leq\dim\mathcal{P}_{k}(M_{i},g)\}$ is open in the $C^{\infty}$ topology.
+Real-analytic metrics are dense in this open set by Greene–Jacobowitz.
+For such a sequence we have $\dim\mathcal{K}(g_{k})\leq\dim\mathcal{K}(g)$, and therefore
+$r^{\sharp}(g_{k})\;:=\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}(g_{k})\;\geq\;r^{\sharp}(g).$
+For each fixed $k$, the curvature tensor $R_{\mathrm{off},k}(s)$ of the connection $\nabla^{(s)}=\nabla^{LC}(g_{k})+\frac{s}{2}T$ is real-analytic in $(s,p)$.
+On the dense open set $U$ furnished by Lemma 5.1 we have $\operatorname{rank}(A_{1,k})=r^{\sharp}(g_{k})$, where $A_{1,k}=\frac{1}{2}(\nabla^{LC}(g_{k})T)_{\mathrm{off}}$.
+Hence there exists an $r^{\sharp}(g_{k})\times r^{\sharp}(g_{k})$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ with $M_{k}(s)=s^{r^{\sharp}(g_{k})}D_{k}(s)$ and $D_{k}(0)\neq 0$.
+By the Identity Principle for Analytic Functions, $D_{k}(s)$ does not vanish identically and its zero set is discrete; consequently, for each $k$ there exists an open–dense subset $U_{k}\subseteq U$ such that
+$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k})\qquad\forall p\in U_{k}.$
+Let $W:=\bigcap_{k=1}^{\infty}U_{k}$.
+Since $M$ is compact, it carries a complete Riemannian metric; in a complete metric space, a countable intersection of open dense subsets is dense.
+Thus $W$ is dense in $U$, hence dense in $M$.
+For every $p\in W$ and every $k$ we have
+$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k}).$
+The rank of a continuous family of linear maps is lower semicontinuous with respect to uniform convergence [9, Warner].
+Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, for every $p\in W$ we obtain
+$\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;\liminf_{k\to\infty}\operatorname{rank}R_{\mathrm{off},k}(p)\;\geq\;\liminf_{k\to\infty}r^{\sharp}(g_{k})\;\geq\;r^{\sharp}(g).$
+The set $\{p\in M\mid\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)\}$ is open (by lower semicontinuity) and contains the dense subset $W$; therefore it is all of $M$.
+Thus $\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)$ for every $p\in M$.
+Finally, by Definition 3.2 the off‑diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ is the Lie subalgebra generated by the off‑diagonal curvature operators and their covariant derivatives at $p$.
+In particular, it contains the linear span of the off‑diagonal curvature operators $R_{\mathrm{off}}(p)$.
+Therefore,
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;r^{\sharp}(g)\qquad\forall p\in M,$
+which is the desired bound.
+The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$.
 ∎
-Remark 5.3 (Necessity of the Non-Parallel Torsion Condition).
-The proof of Theorem 5.1 relies fundamentally on the non-vanishing of the linear term $R^{C}_{\mathrm{off}}\simeq\frac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$, which is established via the $s\to 0$ limit of the one-parameter family of connections $\nabla^{(s)}$.
-This dependence on the linear term necessitates the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof to be valid.
-To illustrate the necessity of this condition, we consider a case where $\nabla^{\mathrm{LC}}T=0$.
-If $\nabla^{\mathrm{LC}}T=0$, the curvature formula simplifies to $R^{C}=R^{\mathrm{LC}}+\frac{1}{4}Q_{T}$, where $Q_{T}$ is the term quadratic in the torsion $T$.
-Since $R^{\mathrm{LC}}$ is block-diagonal for a product metric, the off-diagonal curvature is solely determined by $Q_{T}^{\mathrm{off}}$.
-The proof of Theorem 5.1 provides no guarantee that the quadratic term $Q_{T}^{\mathrm{off}}$ alone is sufficient to generate $r^{\sharp}$ independent directions.
-Consider the flat torus $M=T^{3}=T^{2}\times S^{1}$ endowed with the flat product metric $g$.
-Let $T$ be the constant volume form $T=dx\wedge dy\wedge dz$.
-This form is harmonic, non-trivial, and its mixed rank is $\text{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1$.
-Since $T$ is constant, we have $\nabla^{\mathrm{LC}}T=0$.
-For this specific example, the curvature $R^{C}$ vanishes identically, leading to $\dim(\text{hol}_{p}^{\mathrm{off}}(\nabla^{C}))=0$.
-The lower bound is calculated as $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2}))=1-1=0$.
-The resulting inequality $0\geq 0$ is satisfied, but the bound is trivial.
-This example demonstrates that when $\nabla^{\mathrm{LC}}T=0$, the topological lower bound $r^{\sharp}$ can be reduced to zero, even for a non-trivial cohomology class ($\omega]\neq 0$).
-The failure of the bound to provide a non-trivial constraint confirms that the contribution of the quadratic term $Q_{T}^{\mathrm{off}}$ alone is insufficient to guarantee the non-triviality of the topological lower bound, necessitating the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof of Theorem 5.1 to hold and to yield a meaningful geometric result.
-[^1] The bound is recovered if we consider a non-parallel torsion $T=\omega_{h}+\varepsilon d(\sin(x)dy\wedge dz)$, which maintains the same cohomology class $\omega_{h}]$ but ensures $\nabla^{\mathrm{LC}}T\neq 0$.
+Remark 5.3 (Necessity of the Non‑Parallel Torsion Condition).
+Theorem 5.2 yields a non‑trivial lower bound only when the linear term
+$\tfrac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$
+is non‑zero. If instead $\nabla^{\mathrm{LC}}T=0$, the curvature reduces to
+$R^{C}=R^{\mathrm{LC}}+\tfrac{1}{4}Q_{T},$
+and there is no guarantee that the purely quadratic off‑diagonal part $(Q_{T})_{\mathrm{off}}$ generates $r^{\sharp}$ independent directions.
+Consider for instance the flat torus $M=T^{2}\times S^{1}$ with the constant volume form $T=dx\wedge dy\wedge dz$, which is harmonic and satisfies
+$\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1.$
+Since $T$ is parallel, $\nabla^{\mathrm{LC}}T=0$; hence every simple tensor in any minimal harmonic decomposition is completely parallel. In the notation of Theorem 5.2 we have
+$\mathcal{V}_{2,1}=\operatorname{span}\{dx\wedge dy\otimes dz\},\qquad\mathcal{P}_{2}(T^{2})=\operatorname{span}\{dx\wedge dy\},\qquad\mathcal{P}_{1}(S^{1})=\operatorname{span}\{dz\},$
+so that
+$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1})\bigr)=\mathcal{V}_{2,1},\qquad\ker\widetilde{\Psi}_{p}=0.$
+Thus $\dim\mathcal{K}=1$ and consequently
+$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim\mathcal{K}=1-1=0.$
+A direct computation gives $R^{C}=0$, hence
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})=0,\qquad r^{\sharp}=0.$
+The bound $0\geq 0$ is saturated but trivial, illustrating that without the non‑parallel torsion condition the theorem cannot force irreducibility.
 Remark 5.4 (Geometric Significance of Invariance).
-The invariance of the lower bound under $g_{1}$-deformations that preserve $\mathcal{P}_{2}$ is non-trivial and geometrically significant on manifolds $M_{1}$ with special holonomy (e.g., Calabi-Yau or flat tori).
-In these cases, the space of parallel forms $\mathcal{P}_{2}$ is non-zero, and the existence of non-homothetic deformations preserving $\mathcal{P}_{2}$ is guaranteed by known results in differential geometry (e.g., deformation theory of special holonomy metrics).
-This highlights the topological nature of the bound, which is stable under a non-trivial class of metric variations.
-Remark 5.5 (Forced irreducibility and quantified off-diagonal holonomy).
-Since $\nabla^{C}$ is metric, any non-trivial off-diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$.
-Theorem 5.2 quantifies the “forced irreducibility” principle in terms of the mixed tensor rank, modulo the parallel subspace $\mathcal{P}_{2}\otimes H^{1}(M_{2})$.
-Specifically, the dimension of the off-diagonal holonomy subspace is bounded below by
-$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
-where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$.
-Irreducibility is forced if and only if $r^{\sharp}>0$, i.e.
-when the mixed tensor rank exceeds the dimension of the parallel subspace.
-Remark 5.6 (Analyticity as a proof device).
-The use of real-analytic metrics in the proof of Theorem 5.2 is purely instrumental: analytic approximation (Greene–Jacobowitz) and lower-semicontinuity of matrix rank extend the bound to every $C^{\infty}$ product metric without additional assumptions.
-6. Examples and Sharpness
-The case $M=S^{2}\times\Sigma_{g}$
-Let $\Sigma_{g}$ be a closed oriented surface of genus $g\geq 1$. By Künneth,
-$H^{3}(S^{2}\times\Sigma_{g};\mathbb{R})\cong H^{2}(S^{2})\otimes H^{1}(\Sigma_{g}),$ (6.1)
-since $H^{1}(S^{2})=0$ and $H^{2}(\Sigma_{g})\cong\mathbb{R}$.
-The mixed component is therefore isomorphic to $\mathbb{R}^{2g}$.
-Choosing a harmonic basis $\{\alpha_{j}\}_{j=1}^{2g}$ for $H^{1}(\Sigma_{g})$ and the volume form $\mathrm{vol}_{S^{2}}$ on $S^{2}$, any mixed class has a harmonic representative of the form
-$\omega_{h}=\sum_{j=1}^{2g}c_{j}\,\mathrm{vol}_{S^{2}}\wedge\alpha_{j}.$ (6.2)
-Thus $\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)=2g$.
-For each $j$, select $X,Y\in V_{1}$ so that $\mathrm{vol}_{S^{2}}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\alpha_{j}(Z)\neq 0$; the corresponding curvature endomorphisms produce $2g$ independent off-diagonal directions.
+The invariance of the lower bound $r^{\sharp}$ is a non‑trivial and geometrically significant property. The bound depends only on two pieces of data:
+  • the mixed class $\omega]_{\mathrm{mixed}}$,
+  • the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics.
+Explicitly,
+$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
+where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (resp.
+$\mathcal{V}_{1,2}$) denotes the linear span of the harmonic representatives of the $(2,1)$- (resp.
+$(1,2)$-) Künneth component of $\omega]$.
+Consequently, $r^{\sharp}$ is invariant under any deformation of the product metric $g=g_{1}\oplus g_{2}$ that preserves the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ ($k=1,2$).
+Such deformations include, for instance, any variation of $g_{i}$ inside the connected component of the moduli space of metrics with a fixed set of parallel forms.
+This stability is especially relevant when at least one factor $M_{i}$ possesses special holonomy (e.g., Calabi–Yau manifolds or flat tori), where the spaces $\mathcal{P}^{k}(M_{i})$ are non‑trivial and the metric $g_{i}$ admits large families of deformations that leave $\mathcal{P}^{k}(M_{i})$ unchanged.
+Thus, $r^{\sharp}$ quantifies the part of the topological “mixing strength” of $\omega]$ that cannot be “absorbed” by metric deformations respecting the parallel‑form strata; it captures a genuinely topological obstruction to reducible holonomy.
+Remark 5.5 (Forced irreducibility and quantified off‑diagonal holonomy).
+Since $\nabla^{C}$ is metric, any non‑trivial off‑diagonal holonomy implies that $\mathfrak{hol}_{p}(\nabla^{C})$ acts irreducibly on $T_{p}M$.
+Theorem 5.2 quantifies this “forced irreducibility” principle in terms of the mixed tensor rank, modulo the obstruction space
+$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},$
+where, in the notation of the theorem,
+$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
+Specifically, the dimension of the off‑diagonal holonomy subspace satisfies
+$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
+and irreducibility is forced precisely when the right‑hand side is positive, i.e. when
+$\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}.$
+In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
+Remark 5.6 (Analyticity as a tool of proof).
+The use of real‑analytic metrics in the proof of Theorem 5.2 is purely instrumental.
+By the density theorem of Greene–Jacobowitz [4] and the upper semicontinuity of the dimension of the space of parallel forms, one can approximate any smooth product metric $g=g_{1}\oplus g_{2}$ in the $C^{\infty}$-topology by a sequence of real‑analytic product metrics $\{g^{(k)}\}$ satisfying $\dim\mathcal{P}_{\ell}(M_{i},g^{(k)})\leq\dim\mathcal{P}_{\ell}(M_{i},g)$ for $\ell=1,2$.
+For each analytic metric $g^{(k)}$ the curvature tensor $R_{\mathrm{off},k}(s)$ depends analytically on the deformation parameter $s$ and on the point $p\in M$; consequently the rank of the off‑diagonal curvature is lower‑semicontinuous and can be controlled by the behaviour of analytic minors (Identity Principle).
+Passing to the limit $k\to\infty$ via the lower‑semicontinuity of matrix rank with respect to uniform convergence [9] yields the desired bound for the original $C^{\infty}$ metric $g$.
+Thus the analytic approximation does not impose any restriction on the result—it merely provides a convenient technical tool for extending the estimate from a generic dense subset to the whole manifold.
+6. Illustrative Examples
+The case $\Sigma_{g}\times T^{2}$ (Example of a non-trivial lower bound)
+Fix $g\geq 2$ and endow $\Sigma_{g}$ with any hyperbolic metric and $T^{2}$ with the flat metric. Harmonic spaces:
+$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\qquad H^{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\},$
+and choose a non-zero harmonic $1$-form $\alpha\in H^{1}(\Sigma_{g})$ (such an $\alpha$ exists because $b_{1}(\Sigma_{g})=2g>0$).
+For the hyperbolic metric on $\Sigma_{g}$ we have
+$\mathcal{P}_{1}(\Sigma_{g})=0,\qquad\mathcal{P}_{2}(\Sigma_{g})=\operatorname{span}\{\operatorname{vol}_{\Sigma}\}\cong\mathbb{R},$
+since the volume form is harmonic and parallel.
+For the flat metric on $T^{2}$ all harmonic forms are parallel, hence
+$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
+Consider the mixed cohomology class
+$\omega]=\underbrace{\operatorname{vol}_{\Sigma}\otimes dx}_{\in H^{2}(\Sigma_{g})\otimes H^{1}(T^{2})}\;+\;\underbrace{\alpha\otimes\operatorname{vol}_{T^{2}}}_{\in H^{1}(\Sigma_{g})\otimes H^{2}(T^{2})}\;\in\;H^{3}(\Sigma_{g}\times T^{2}).$
+Both Künneth components are already simple tensors, therefore their minimal tensor ranks are
+$r_{2,1}=1,\qquad r_{1,2}=1,$
+and the mixed rank equals
+$\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)=r_{2,1}+r_{1,2}=2.$
+The obstruction kernels are computed as follows. For the $(2,1)$-part $\pi_{2,1}([\omega])=\operatorname{vol}_{\Sigma}\otimes dx$,
+$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)=\operatorname{span}\{\operatorname{vol}_{\Sigma}\otimes dx\},$
+because $\operatorname{vol}_{\Sigma}\in\mathcal{P}_{2}(\Sigma_{g})$ and $dx\in\mathcal{P}_{1}(T^{2})$. Hence $\dim\ker\Psi_{p}=1$.
+For the $(1,2)$-part $\pi_{1,2}([\omega])=\alpha\otimes\operatorname{vol}_{T^{2}}$,
+$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)=\{0\},$
+because $\mathcal{P}_{1}(M_{1})=\{0\}$ and $\mathcal{P}_{2}(M_{2})=\operatorname{span}\{\operatorname{vol}_{T^{2}}\}$, hence $\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})=\{0\}$.
+Hence $\dim\ker\widetilde{\Psi}_{p}=0$.
+Thus
+$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\dim\mathcal{K}=1+0=1,$
+and the corrected rank is
+$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}=2-1=1.$
+Theorem 5.2 therefore yields the non‑trivial lower bound
+$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;1.$
+In particular, any cohomologically calibrated metric connection with non‑parallel torsion representing $\omega]$ necessarily has irreducible holonomy across the product splitting $\Sigma_{g}\times T^{2}$.
+The case $T^{2}\times T^{2}$ (Example of a trivial lower bound)
+Let $M=T^{2}\times T^{2}$ endowed with the flat product metric $g=g_{1}\oplus g_{2}$. Fix harmonic bases
+$H^{1}(T^{2})=\operatorname{span}\{dx,dy\},\quad H^{2}(T^{2})=\operatorname{span}\{vol\}.$
+For the flat metric we have
+$\mathcal{P}_{1}(T^{2})=H^{1}(T^{2})\cong\mathbb{R}^{2},\qquad\mathcal{P}_{2}(T^{2})=H^{2}(T^{2})\cong\mathbb{R}.$
+The mixed Künneth components are
+$H^{2}(T^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2},\quad H^{1}(T^{2})\otimes H^{2}(T^{2})\cong\mathbb{R}^{2}.$
+Consider the single cohomology class
+$\omega]=\underbrace{vol\otimes dx}_{(2,1)}+\underbrace{dy\otimes vol}_{(1,2)}.$
+Its minimal tensor-rank decomposition is
+$\pi_{2,1}([\omega])=vol\otimes dx\quad(\text{rank }1),\qquad\pi_{1,2}([\omega])=dy\otimes vol\quad(\text{rank }1).$
 Hence
-$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2g-\dim(\mathcal{P}_{2}\otimes H^{1}(\Sigma_{g})),$ (6.3)
-where $\mathcal{P}_{2}\subset H^{2}(S^{2};\mathbb{R})$ is the subspace of harmonic parallel $2$-forms on $S^{2}$ (here $\mathcal{P}_{2}=\{0\}$, so the bound is $2g$).
-The case $M=S^{2}\times T^{2}$
-Here $H^{3}(M)\cong H^{2}(S^{2})\otimes H^{1}(T^{2})\cong\mathbb{R}^{2}$.
-Writing $\omega_{h}=c_{1}\,\mathrm{vol}_{S^{2}}\wedge dx+c_{2}\,\mathrm{vol}_{S^{2}}\wedge dy$ with $dx,dy$ the harmonic $1$-forms on the torus, one obtains two independent mixed components, hence
-$\dim\big(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\big)\ \geq\ 2-\dim(\mathcal{P}_{2}\otimes H^{1}(T^{2})),$ (6.4)
-where $\mathcal{P}_{2}=\{0\}$ on $S^{2}$, so the bound is $2$.
+$\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1+1=2.$
+Since $vol\in\mathcal{P}_{2}$ and $dx,dy\in\mathcal{P}_{1}$, both simple tensors lie in the obstruction spaces
+$\ker\Psi_{p}=\operatorname{span}\{vol\otimes dx\},\qquad\ker\tilde{\Psi}_{p}=\operatorname{span}\{dy\otimes vol\},$
+so $\dim\mathcal{K}=2$. Theorem 5.2 gives
+$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
+7. Significance and Applications to Physics and STEM Disciplines
+Connections with torsion and product structures appear naturally in a wide range of mathematical and physical models.
+In string theory, supergravity, and related frameworks, metric connections with totally skew-symmetric torsion play a central role in flux compactifications, generalized geometries, and effective field theories on product manifolds such as $M_{4}\times X$.
+In these settings, the holonomy of the canonical connection governs the coupling between the external (space-time) and internal (compactified) factors, and irreducibility prevents the decoupling of physical degrees of freedom.
+Our lower bound provides an efficient mechanism for establishing such irreducibility directly from global invariants - particularly the mixed tensor rank and the corrected quantity $r^{\sharp}$ - without computing curvature tensors or higher covariant derivatives.
+This is especially relevant in high-dimensional backgrounds where explicit curvature calculations are notoriously difficult or mathematically intractable.
+A similar phenomenon arises in gauge theory, condensed matter, and materials science, where product manifolds model layered, composite, or multi-component systems.
+The holonomy algebra encodes the geometric interactions between these components; detecting whether such interactions are essential (irreducible) or merely apparent becomes a fundamental structural question.
+The criterion developed in this paper identifies irreducibility based solely on the existence and non-parallelity of mixed differential forms, providing a geometric diagnostic that is robust under perturbations of the underlying metric and torsion.
+Applications also extend to geometric analysis and computational modeling.
+In multi-scale PDE systems defined on product domain (such as transport-diffusion equations, elasticity models, or geometric flows) the holonomy of a natural geometric connection determines how different subsystems interact and whether the governing equations admit a decoupled formulation.
+Our lower bound offers a rigorous obstruction to such decoupling, enabling researchers to understand when full coupling is intrinsic to the geometry.
+This is particularly useful in numerical simulations, where direct computation of curvatures is expensive and unstable, while global invariants such as parallel-form spaces are straightforward to detect algorithmically.
+Overall, the ability to certify irreducible holonomy through topological or geometric invariants provides a powerful tool across physics, geometry, and applied sciences.
+It eliminates the need for explicit curvature computations and transforms a classically analytic problem into a global geometric one, opening the door to applications in any setting where torsion, product structures, and holonomy play a foundational role.
 8. Concluding Remarks and Perspectives
-We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with totally skew torsion calibrated by a mixed cohomology class, the off-diagonal holonomy dimension admits a quantified lower bound
-$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2})),$
-where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ is the space of harmonic parallel $2$-forms on $M_{1}$.
+We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with 3-form torsion representing a mixed cohomology class via its harmonic projection $\omega_{h}$, the off-diagonal holonomy dimension admits a quantified lower bound
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K},$
+where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\tilde{\Psi}_{p}$ is the sum of the kernels of the linear maps $\Psi_{p},\tilde{\Psi}_{p}$ that encode the obstruction to off-diagonal curvature.
 The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds.
-This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace.
-The examples above demonstrate sharpness in natural settings where the torsion equals the harmonic representative and the parallel subspace has minimal dimension.
-It is natural to conjecture that equality holds whenever the torsion is purely harmonic, whereas strict inequality may occur when the coexact or exact parts of $T$ generate additional independent off-diagonal curvature directions or when the parallel subspace has maximal dimension.
-A systematic classification of such cases is an interesting direction for future work.
+This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace contained within the mixed class.
+The examples above illustrate the computation of $r^{\sharp}$ in simple geometric settings.
+When the torsion is purely harmonic and the parallel-form spaces are minimal, the bound is often saturated in explicit constructions (e.g., on flat tori with parallel torsion).
+It is natural to conjecture that equality holds generically for purely harmonic torsion, while strict inequality may occur when the coexact or exact parts of $T$ generate additional off-diagonal curvature directions, or when the parallel-form spaces are larger.
+A systematic classification of the cases where the bound is sharp remains an interesting direction for future work.
 A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].

### v10

@@
 Abstract.
-We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of metric connections with totally skew torsion on product manifolds.
-For any such connection whose non‑parallel torsion 33‑form represents a mixed de Rham class via its harmonic projection, we prove that the dimension of the off‑diagonal holonomy subspace satisfies the lower bound
-$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
-where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the linear obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$.
-These kernels are defined intrinsically as the intersections
-$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
-where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (respectively $\mathcal{V}_{1,2}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$) denotes the linear span of the harmonic representatives of the $(2,1)$- (respectively $(1,2)$-) Künneth component of $\omega]$.
-The subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are independent of the choice of a minimal harmonic decomposition.
-Consequently, the integer $\dim\mathcal{K}$ depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics; hence the bound is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
-This yields a computable, metric‑invariant obstruction to reducible holonomy and quantifies the topological “mixing strength” that forces irreducibility across the product splitting.
+We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting and the mixed class has tensor rank $1$.
+For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
+with $\mathcal{K}$ an intrinsically defined obstruction space.
+The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides a sharp obstruction to reducible holonomy.
+Counterexamples show the hypothesis is optimal.
 Keywords: Connections with Torsion, Product Manifolds, Holonomy Algebra, De Rham Cohomology, Hodge Theory, Künneth Decomposition, Curvature Tensor
@@
 We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
-More precisely, we prove
+A subtle but crucial point is that the subspaces spanned by a minimal harmonic decomposition are uniquely determined only under topological restrictions on the factors.
+Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e.:
+  • (A)
+    both factors are compact oriented surfaces;  or
+  • (B)
+    $\dim M=4$ with a $(3,1)$ splitting and the mixed class has total tensor rank $1$.
+For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies
 $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
-where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the natural linear obstruction maps. To define these kernels, let
-$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
-where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp.
-$\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp.
-$(1,2)$-) Künneth component of $\omega]$.
-Because any two minimal harmonic decompositions span the same subspaces $U\subset\mathcal{H}^{2}(M_{1})$ and $V\subset\mathcal{H}^{1}(M_{2})$, the spaces $\mathcal{V}_{2,1}=U\otimes V$ and $\mathcal{V}_{1,2}$ are well‑defined, i.e., independent of the chosen decomposition.
-Then
-$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
-The integer $\dim\mathcal{K}$ measures precisely the part of the mixed class that is absorbed by the parallel‑form strata and therefore does not contribute to off‑diagonal curvature.
-The quantity $r^{\sharp}$ is metric‑invariant under deformations that preserve the parallel‑form strata: it depends only on $\omega]_{\mathrm{mixed}}$ and on the subspaces $\mathcal{P}_{k}(M_{i})$.
-Consequently, each independent mixed Künneth tensor guarantees a linearly independent off‑diagonal holonomy direction, modulo the parallel forms captured by $\mathcal{K}$.
-This yields a computable, purely topological criterion for forced irreducibility of the holonomy representation across the product splitting: irreducibility is forced whenever $r^{\sharp}>0$, i.e.
-whenever $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}$.
+where the obstruction space $\mathcal{K}$ is intrinsically defined via intersections of the mixed factor spaces $\mathcal{V}_{2,1}$, $\mathcal{V}_{1,2}$ with the parallel‑form strata $\mathcal{P}_{k}(M_{i})$.
+The bound is invariant under metric deformations preserving these strata.
+When $r^{\sharp}>0$ the holonomy representation is forced to be irreducible across the product splitting.
+The examples in Section 6 illustrate both non‑trivial ($\Sigma_{g}\times T^{2}$) and trivial ($T^{2}\times T^{2}$) cases, and Remark 2.4 shows by explicit counterexample that without the admissibility hypothesis the bound fails to be well‑defined.
 2. Cohomologically Calibrated Metric Connections with Skew Torsion
@@
 For a thorough treatment of Hodge theory and cohomology on complex manifolds, see [5].
+2.1. Intrinsic Definition of Mixed Cohomology Subspaces
+We now address a subtle but crucial point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general.
+To obtain a well–defined geometric invariant we must impose conditions on the topology of the factors.
+Definition 2.2 (Admissible Product Manifolds).
+A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies one of the following:
+  • (A)
+    both factors $M_{1},M_{2}$ are compact oriented surfaces;
+  • (B)
+    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
+Lemma 2.3 (Uniqueness of Mixed Subspaces).
+Let $M$ be an admissible product manifold and $\omega]\in H^{3}(M;\mathbb{R})$ a non–trivial mixed class. Then the subspaces
+$\mathcal{V}_{2,1}$ $\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$ (2.7)
+$\mathcal{V}_{1,2}$ $\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$ (2.8)
+spanned by any minimal harmonic decomposition of the mixed Künneth components of $\omega]$ are intrinsically defined—i.e.
+they depend only on $\omega]$ and not on the particular minimal decomposition chosen.
+Proof.
+We treat the two admissible cases separately.
+Case (A): $\dim M_{i}=2$. For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
+$\mathcal{H}^{2}(F)=\mathbb{R}\cdot\operatorname{vol}_{F}.$
+Consequently any non–zero $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank one and can be written uniquely as
+$T=\operatorname{vol}_{M_{1}}\otimes\beta,\qquad\beta\in\mathcal{H}^{1}(M_{2}).$
+Indeed, if $\operatorname{vol}_{M_{1}}\otimes\beta=\operatorname{vol}_{M_{1}}\otimes\beta^{\prime}$, then $\operatorname{vol}_{M_{1}}\otimes(\beta-\beta^{\prime})=0$ and, because $\operatorname{vol}_{M_{1}}$ is non–zero as an element of $\mathcal{H}^{2}(M_{1})$, we must have $\beta=\beta^{\prime}$.
+Thus the one–dimensional subspace $\operatorname{span}\{\beta\}$ is uniquely determined by $T$, and
+$\mathcal{V}_{2,1}=\mathcal{H}^{2}(M_{1})\otimes\operatorname{span}\{\beta\}$
+is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$.
+Case (B): $(3,1)$ or $(1,3)$ with $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
+Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric).
+Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$.
+Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$.
+Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
+$T=\alpha\otimes\gamma,\qquad\alpha\in\mathcal{H}^{2}(M_{1}).$
+Uniqueness follows as in Case (A) because $\gamma$ is non–zero.
+Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ depends only on $\omega]$.
+The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined.
+∎
+Remark 2.4 (Failure of invariance in general).
+Invariance fails without the admissibility hypothesis. For instance, on $M=T^{3}\times T^{3}$ we have $b_{2}(T^{3})=3$ and $b_{1}(T^{3})=3$. The mixed class
+$\omega]=(dx\wedge dy)\otimes dz+(dx\wedge dz)\otimes dy\in H^{2}(T^{3})\otimes H^{1}(T^{3})$
+has rank $2$. It admits two distinct minimal decompositions
+$\omega_{h}^{2,1}$ $=(dx\wedge dy)\otimes dz+(dx\wedge dz)\otimes dy$
+$=\tfrac{1}{2}\big[(dx\wedge dy+dx\wedge dz)\otimes(dz+dy)$
+$\qquad\quad+(dx\wedge dy-dx\wedge dz)\otimes(dz-dy)\big],$
+which generate different subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{2,1}^{\prime}$.
+Consequently the integer $\dim\big(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}))\big)$ depends on the choice, contradicting the metric–invariance.
+2.2. Mixed obstruction spaces
+Having established the intrinsic nature of the mixed factor spaces in Lemma 2.3 for admissible product manifolds (Definition 2.2), we now define the obstruction kernels.
+Definition 2.5 (Mixed obstruction spaces).
+For an admissible product manifold $M$ and a mixed class $\omega]$ we define
+$\mathcal{V}_{2,1}$ $:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$
+$\mathcal{V}_{1,2}$ $:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
+where the spans are taken over any minimal harmonic decomposition of the mixed Künneth components of $\omega]$; by Lemma 2.3 these spans are independent of the particular choice.
+The mixed obstruction spaces are the intersections
+$\ker\Psi_{p}$ $:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
+$\ker\widetilde{\Psi}_{p}$ $:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
+and we set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
 3. Off-Diagonal Curvature and Holonomy Complexity
@@
 $R_{\mathrm{off}}(X,Y)Z:=P_{1}R(X,Y)(P_{2}Z)+P_{2}R(X,Y)(P_{1}Z).$ (3.1)
-Definition 3.2 (Off-diagonal holonomy subspace).
-Let $(M,g)$ be a Riemannian manifold and $\nabla^{C}$ a metric connection on $TM$ with totally skew–symmetric torsion.
-For a point $p\in M$, we define the off–diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ as the Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ generated by all off–diagonal components of the curvature tensor and its covariant derivatives at $p$, namely
-$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C}):=\mathrm{Lie}\,\Big\langle\big(\nabla^{k}R^{C}\big)_{p}(X_{1},\ldots,X_{k};Y,Z)^{\mathrm{off}}\;\Big|\;k\geq 0,\;X_{i},Y,Z\in T_{p}M\Big\rangle,$
-where $(\cdot)^{\mathrm{off}}$ denotes the off–diagonal projection with respect to the product splitting $T_{p}M=V_{1}\oplus V_{2}$.
+Definition 3.2 (Off-diagonal holonomy algebra).
+Let $(M,g)$ be a Riemannian product manifold with orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$, and let $\nabla^{C}$ be a metric connection on $TM$ with totally skew-symmetric torsion.
+Let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
+$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z),$
+where $P_{i}:T_{p}M\to V_{i}$ are the orthogonal projections.
+Let $\mathcal{G}_{p}:=\{\nabla^{k}R^{C}|_{p}(X_{1},\dots,X_{k};Y,Z)\mid k\geq 0,X_{i},Y,Z\in T_{p}M\}$ be the set of Ambrose–Singer curvature generators of $\mathfrak{hol}_{p}(\nabla^{C})$.
+Define the off-diagonal curvature span
+$\mathcal{S}_{p}^{\mathrm{off}}:=\operatorname{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(A)\mid A\in\mathcal{G}_{p}\}\subset\mathfrak{so}(T_{p}M).$
+The off-diagonal holonomy algebra is the Lie subalgebra obtained by intersecting $\mathfrak{hol}_{p}(\nabla^{C})$ with the Lie algebra generated by $\mathcal{S}_{p}^{\mathrm{off}}$ in $\mathfrak{so}(T_{p}M)$:
+$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\bigl\langle\mathcal{S}_{p}^{\mathrm{off}}\bigr\rangle_{\mathrm{Lie}}\cap\mathfrak{hol}_{p}(\nabla^{C})\subset\mathfrak{hol}_{p}(\nabla^{C}).$
+Here $\langle\cdot\rangle_{\mathrm{Lie}}$ denotes the Lie algebra generated in $\mathfrak{so}(T_{p}M)$, and the intersection ensures that $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is canonically a subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$.
+Equivalently, it is the smallest Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ whose off-diagonal components contain $\mathcal{S}_{p}^{\mathrm{off}}$.
 By the Ambrose–Singer theorem [2], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor.
@@
 5. Main Result and Proof
-Lemma 5.1 (Generic rank of covariant derivatives).
-Let $(M_{1},g_{1})$ be a compact and oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms.
-Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$.
-Then the set
-$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\operatorname{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
-is open and dense in $M_{1}$.
+Lemma 5.1 (Existence of points of maximal rank for covariant derivatives).
+Let $(M_{1},g_{1})$ be a compact oriented Riemannian manifold and let $\mathcal{H}^{2}(M_{1},g_{1})$ denote the space of harmonic $2$-forms.
+Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$.
+Then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
+$\bigl(\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}\bigr)$
+are pointwise linearly independent; i.e.
+$\mathcal{U}=\Bigl\{p\in M_{1}\;\Big|\;\operatorname{rank}\bigl[(\nabla^{g_{1}}\alpha_{i})(p)\bigr]_{i=s+1,\dots,r}=r-s\Bigr\}\neq\emptyset.$
 Proof.
-Fix smooth local frames of $\Lambda^{2}T^{*}M_{1}$ and $T^{*}M_{1}$ over an open coordinate ball $B\subset M_{1}$.
-In these frames each $(\nabla^{g_{1}}\alpha_{i})(p_{1})$ becomes a smooth $\mathbb{R}$-valued matrix $A_{i}(p_{1})$ of size $\dim M_{1}\times\binom{\dim M_{1}}{2}$.
-Form the $(r_{2,1}-s)\times\bigl[\dim M_{1}\cdot\binom{\dim M_{1}}{2}\bigr$] matrix
-$\mathbf{A}(p_{1})=\bigl[A_{s+1}(p_{1})^{\top}\!,\dots,A_{r_{2,1}}(p_{1})^{\top}\bigr]^{\top}.$
-The condition $\operatorname{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non‑vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$.
-Each minor is a smooth function on $B$; hence its non‑zero locus is open in $B$.
-Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open.
-Now suppose, for a contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non‑empty open set $V\subset M_{1}$ on which
-$\operatorname{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
-Since the rank is lower‑semicontinuous, after possibly shrinking $V$ we may assume that the rank is constant, say $\rho<r_{2,1}-s$, on $V$.
-Consequently, there exists a smooth map $\lambda:V\to\mathbb{R}^{r_{2,1}-s}\setminus\{0\}$ such that
-$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{1})(\nabla^{g_{1}}\alpha_{i})(p_{1})=0\qquad\forall p_{1}\in V.$ (5.1)
-Now we employ analytic approximation.
-By the theorem of Greene–Jacobowitz [4], we can approximate the metric $g_{1}$ in the $C^{\infty}$ topology by a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}$.
-For each $k$, denote by $\alpha_{i}^{(k)}$ the $g_{1}^{(k)}$-harmonic representative of the cohomology class $\alpha_{i}]$.
-The forms $\alpha_{i}^{(k)}$ converge to $\alpha_{i}$ in $C^{\infty}$ as $k\to\infty$.
-Because $\lambda$ is smooth and non‑zero on $V$, we may choose a point $p_{0}\in V$ and a subsequence (still denoted by $k$) such that $\lambda(p_{0})\neq 0$ and
-$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})(\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)})(p_{0})=0\qquad\text{for all }k.$ (5.2)
-Indeed, (5.1) passes to the limit after extracting a suitable subsequence, since all objects depend smoothly on the metric.
-For each fixed $k$, the metric $g_{1}^{(k)}$ is real‑analytic. Define
-$\eta^{(k)}:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}^{(k)}\in\mathcal{H}^{2}(M_{1},g_{1}^{(k)}).$
-Equation (5.2) states that $\nabla^{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$.
-Because $g_{1}^{(k)}$ is analytic, the coefficients of the Hodge–Laplacian are analytic, and Aronszajn’s unique-continuation theorem [1] applies.
-Since $\eta^{(k)}$ is harmonic and $g_{1}^{(k)}$ is real-analytic, $\eta^{(k)}$ is also real-analytic.
-The condition $\nabla_{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$ implies that $\eta^{(k)}$ is parallel in a neighborhood of $p_{0}$.
-By the unique-continuation principle for parallel forms on real-analytic manifolds (a consequence of Aronszajn’s theorem applied to the analytic coefficients), $\eta^{(k)}$ must be parallel on the entire connected component.
-Passing to the limit $k\to\infty$, we obtain that
-$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}$
-is parallel with respect to the original metric $g_{1}$ on the component containing $p_{0}$.
-Thus $\eta\in\mathcal{P}_{2}(M_{1})$, contradicting the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}(M_{1})$.
-Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
-The same reasoning applies verbatim to $\mathcal{H}^{k}(M_{i})$ for any $k$ and for both factors $M_{1}$ and $M_{2}$. ∎
+We proceed in three steps.
+1.
+Reduction to a real‑analytic metric.
+Real‑analytic Riemannian metrics are dense in the space of smooth metrics on a compact manifold; this follows from the theorem of Greene–Jacobowitz on analytic approximation of smooth maps (see [4]).
+Hence we can choose a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}_{k\in\mathbb{N}}$ converging to $g_{1}$ in the $C^{\infty}$ topology.
+The dimension of the space of parallel $2$-forms is upper‑semicontinuous with respect to the metric (see [3]); therefore, after passing to a subsequence, we may assume
+$\dim\mathcal{P}_{2}\bigl(M_{1},g_{1}^{(k)}\bigr)\leq\dim\mathcal{P}_{2}(M_{1},g_{1})=s\qquad\text{for all }k.$
+For each $k$ let $\alpha_{i}^{(k)}\in\mathcal{H}^{2}\bigl(M_{1},g_{1}^{(k)}\bigr)$ be the unique harmonic representative of the cohomology class $\alpha_{i}]$.
+By the stability theorem for harmonic forms under metric deformation (Warner [9], see also Kodaira–Spencer [7]), $\alpha_{i}^{(k)}\to\alpha_{i}$ in $C^{\infty}$.
+Consequently the covariant derivatives converge in $C^{1}$:
+$\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}\longrightarrow\nabla^{g_{1}}\alpha_{i}\qquad(i=s+1,\dots,r).$
+If we prove that for each $k$ there exists a non‑empty open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a non‑empty open set for the original metric $g_{1}$.
+Indeed, linear independence is equivalent to the non‑vanishing of at least one $(r-s)\times(r-s)$ minor of the matrix of components; each minor depends continuously on the metric and the forms, so a minor that is non‑zero on an open set for $g_{1}^{(k)}$ remains non‑zero on an open set for $g_{1}$ when $k$ is sufficiently large.
+Thus it suffices to prove the lemma under the assumption that $g_{1}$ is real‑analytic.
+2.
+The real‑analytic case.
+Assume now that $g_{1}$ is real‑analytic.
+Then every harmonic form is real‑analytic (by elliptic regularity for the Hodge–Laplacian, see [8]), and consequently the covariant derivatives $\nabla^{g_{1}}\alpha_{i}$ are real‑analytic sections of the vector bundle $E:=T^{*}M_{1}\otimes\Lambda^{2}T^{*}M_{1}$.
+Consider the real‑analytic section
+$\Phi:M_{1}\longrightarrow\Lambda^{r-s}E,\qquad\Phi(p):=(\nabla^{g_{1}}\alpha_{s+1})(p)\wedge\cdots\wedge(\nabla^{g_{1}}\alpha_{r})(p).$
+The set $\mathcal{U}:=\{p\in M_{1}\mid\Phi(p)\neq 0\}$ is exactly the set where the covariant derivatives are linearly independent; because $\Phi$ is continuous, $\mathcal{U}$ is open.
+We must show that $\mathcal{U}\neq\emptyset$.
+Suppose, for contradiction, that $\mathcal{U}=\emptyset$; i.e.
+$\Phi(p)=0$ for every $p\in M_{1}$.
+Then at each point the sections $\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}$ are linearly dependent.
+Since the condition of rank $\leq d$ is given by the vanishing of all $(d+1)\times(d+1)$ minors, which are analytic functions, the locus where the rank is constant is an analytic subset.
+Let d<rsd<r-s be the minimal rank and let $p_{0}$ be a point where the rank equals $d$.
+There exists a connected open neighbourhood $V$ of $p_{0}$ on which the rank is constant and equal to $d$.
+On $V$ we can therefore find real‑analytic functions $c_{1},\dots,c_{r-s}:V\to\mathbb{R}$, not all zero, such that
+$\sum_{i=1}^{r-s}c_{i}(p)\bigl(\nabla^{g_{1}}\alpha_{s+i}\bigr)(p)=0\qquad\forall p\in V.$ (5.1)
+Define the analytic $2$-form $\eta:=\sum_{i=1}^{r-s}c_{i}\alpha_{s+i}$ on $V$.
+From (5.1) we have $\nabla^{g_{1}}\eta=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}+\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}$.
+Let $V\subset M_{1}$ be the non–empty open subset fixed above.
+Since the mixed cohomology class on $M_{1}\times M_{2}$ is assumed to be non–trivial, its harmonic representative has a non–vanishing component in $H^{2}(M_{1})\otimes H^{1}(M_{2})$.
+In particular, the projection onto $H^{2}(M_{1})$ is non–zero.
+Hence the harmonic forms $\alpha_{s+1},\dots,\alpha_{r}$ cannot all vanish identically on $M_{1}$.
+Let $W\subset M_{1}$ denote the complement of their common zero set.
+By real–analyticity, $W$ is a non–empty open subset of $M_{1}$.
+Restricting to the non‑empty open set $V^{\prime}:=V\cap W$, we have on $V^{\prime}$
+$\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}=0.$
+Because the $\alpha_{s+i}(p)$ are linearly independent at every $p\in V^{\prime}$, this forces $dc_{i}=0$ on $V^{\prime}$ for each $i$.
+Thus the functions $c_{i}$ are constant on each connected component of $V^{\prime}$.
+By analyticity, they extend to constants on the whole connected open set $V$ (if two analytic functions coincide on an open set, they coincide everywhere on the connected component).
+Consequently, the $c_{i}$ are constant on $V$.
+Thus there exist constants $c_{1},\dots,c_{r-s}\in\mathbb{R}$, not all zero, such that
+$\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=0\qquad\text{on }V.$
+Define the global harmonic $2$-form
+$\zeta:=\sum_{i=1}^{r-s}\bar{c}_{i}\alpha_{s+i}\in\mathcal{H}^{2}(M_{1}).$
+Then $\nabla^{g_{1}}\zeta=0$ on $V$; i.e., $\zeta$ is parallel on the open set $V$.
+Because both $\zeta$ and the condition $\nabla^{g_{1}}\zeta=0$ are real-analytic, the unique-continuation principle [1] implies that $\nabla^{g_{1}}\zeta=0$ on the whole connected component of $M_{1}$ containing $V$, hence on all of $M_{1}$.
+Therefore $\zeta$ is a globally parallel $2$-form.
+But this contradicts the hypothesis that $\alpha_{s+1},\dots,\alpha_{r}$ are linearly independent modulo $\mathcal{P}_{2}(M_{1},g_{1})$.
+Indeed, $\zeta$ is a non-trivial linear combination of these forms that belongs to $\mathcal{P}_{2}(M_{1},g_{1})$.
+Hence our assumption that $\mathcal{U}=\emptyset$ is impossible, and we conclude that $\mathcal{U}$ is a non-empty open subset of $M_{1}$.
+3.
+Returning to the original smooth metric.
+As explained in Step 1, the existence of a non‑empty open set of linear independence for a sequence of real‑analytic approximations $g_{1}^{(k)}$ together with the $C^{1}$ convergence of the covariant derivatives guarantees the existence of a non‑empty open set with the same property for the original smooth metric $g_{1}$.
+This completes the proof.
+∎
 Theorem 5.2 (Topological Lower Bound for Off-Diagonal Curvature).
-Let $M=M_{1}\times M_{2}$ be a compact and oriented product manifold with a product metric $g=g_{1}\oplus g_{2}$, and let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
-Let
-$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
-where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp.
-$\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp.
-$(1,2)$-) Künneth component of $\omega]$.
-Define
-$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
-and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$. Then, for every $p\in M$,
+Let $M=M_{1}\times M_{2}$ be an admissible compact oriented Riemannian product manifold (Definition 2.2) with product metric $g=g_{1}\oplus g_{2}$.
+Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class.
+Define the mixed factor subspaces
+$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),$
+$\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
+where the spans are taken over any minimal harmonic decomposition of the mixed Künneth components of $\omega]$; in the admissible setting these spans are independent of the particular choice (Lemma 2.3).
+Define the obstruction kernels
+$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
+and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
+Then, for every $p\in M$,
 $\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
-The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
+The bound depends only on $\omega]_{\text{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
 Proof.
@@
 Consider the curvature formula (2.5).
-The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources.
-First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$.
-Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off‑diagonal operators.
-Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$.
-Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
+The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts.
+Our aim is to show that the off‑diagonal holonomy algebra contains at least $r^{\sharp}$ linearly independent directions.
+These directions will be produced by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the harmonic part $\omega_{h}$ alone.
+We must verify that the remaining contributions—the part of the linear term coming from $\eta$ and the quadratic term $Q_{T}$—cannot interfere with this lower bound.
+Suppose that on some non‑empty open set $V\subset M$, the derivative $\nabla^{LC}\eta$ cancelled enough off‑diagonal directions from $\nabla^{LC}\omega_{h}$ to reduce the rank of $(\nabla^{LC}T)_{\mathrm{off}}$ below $r^{\sharp}$.
+This would impose linear relations among the covariant derivatives of the harmonic components $\alpha_{i}\wedge\beta_{i}$ on $V$.
+Applying the analytic‑approximation technique employed in Lemma 5.1 (see also Remark 5.6), we may work with real‑analytic metrics and forms.
+The relation then becomes an equality between real‑analytic sections, so by the unique continuation principle for analytic sections (Aronszajn [1]) it extends to the whole connected component containing $V$, hence globally.
+Since the tensors $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})$ are pointwise linearly independent on the open set $\mathcal{U}$ provided by Lemma 5.1, evaluating the global relation at any point of $\mathcal{U}$ forces all coefficients to vanish, contradicting the existence of a non‑trivial cancellation.
+Hence, on a (possibly smaller) non‑empty open set, the linear term $(\nabla^{LC}T)_{\mathrm{off}}$ still exhibits at least $r^{\sharp}$ independent off‑diagonal directions originating from $\omega_{h}$.
+Regarding the quadratic term $Q_{T}$, observe that the $r^{\sharp}$ independent operators produced by the linear term belong to the off‑diagonal holonomy algebra by the Ambrose–Singer theorem [2].
+The quadratic term contributes additional operators to the same algebra.
+Since the dimension of a vector space generated by a set of vectors is at least the rank of any linearly independent subset, adding further generators cannot reduce the dimension below $r^{\sharp}$.
+Consequently, the lower bound obtained from the linear term persists for the full curvature $R^{C}$.
+Thus, to establish $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ it suffices to analyse the linear term generated by the harmonic component $\omega_{h}$.
 To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$.
-By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open‑dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent.
+By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a non‑empty open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent.
 For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
 Define the linear map
-$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}).$ (5.5)
-The map is well‑defined and linear; its image consists precisely of the off‑diagonal curvature operators generated by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the (2,1)(2,1)‑component of $\omega_{h}$.
-If $\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$, then the corresponding off‑diagonal curvature contribution vanishes:
+$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.4)
+where the vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfy $\alpha_{i}(X_{i},Y_{i})=1$ and $\beta_{i}(Z_{i})=1$.
+Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.4.
+Proof of Claim.
+Writing $\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i}$, the product rule gives
+$(\nabla^{LC}\omega_{h})^{\mathrm{off}}=\sum_{i=1}^{r_{2,1}}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i}).$
+For vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$, $Z_{i}\in T_{p_{2}}M_{2}$ with $\alpha_{i}(X_{i},Y_{i})=1$, $\beta_{i}(Z_{i})=1$, the contraction $\iota_{Z_{i}}\iota_{Y_{i}}\iota_{X_{i}}$ extracts precisely the $i$-th component:
+$(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i};Z_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i}).$
+Summing over $i$ with coefficients $c_{i}$ yields that $\Psi_{p}(\xi)=0$ is equivalent to the tensor equation
+$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0,$ (5.5)
+which is manifestly independent of the vectors $\{X_{i},Y_{i},Z_{i}\}$. Hence $\ker\Psi_{p}$ is well-defined. ∎
+Now, although the right-hand side of   5.4 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
+If $\xi\in\ker\Psi_{p}$, then by definition
 $\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
-Using the product rule $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})$ and the fact that $R^{LC}_{\mathrm{off}}=0$, this condition is equivalent to
-$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0.$ (5.6)
-Separating the bidegrees in (5.6) yields the two independent conditions
-$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.7)
-Because the forms $\beta_{i}$ (resp. $\alpha_{i}$) are linearly independent on an open‑dense set, we obtain
-$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}=0$ (5.8)
-on that set. By the unique‑continuation principle for harmonic forms (Lemma 5.1), these equalities extend globally, hence
-$\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.9)
+Separating the bidegrees in   5.5 yields the two independent conditions
+$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\quad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.6)
+Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a non‑empty open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants).
+Denote this open set by $\mathcal{U}_{2}\subset M_{2}$.
+Now fix any point $y_{0}\in\mathcal{U}_{2}$. For this fixed $y_{0}$, the first equation in (5.6) becomes:
+$\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)\Bigr)\wedge\beta_{i}(y_{0})=0\quad\text{for all }x\in M_{1}.$
+Since the covectors $\beta_{1}(y_{0}),\dots,\beta_{r_{2,1}}(y_{0})$ are linearly independent, this forces
+$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)=0\quad\text{for all }x\in M_{1}.$
+Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ where they are pointwise linearly independent.
+Fix $x_{0}\in\mathcal{U}_{1}$. The second equation in (5.6) then gives for all $y\in M_{2}$:
+$\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}(x_{0})\Bigr)\wedge\nabla^{g_{2}}\beta_{i}(y)=0,$
+which implies $\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(y)=0$ for all $y\in M_{2}$.
+Therefore,
+$\eta_{1}:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.7)
 Conversely, if (5.7) holds, then $\nabla^{LC}\bigl(\sum_{i}c_{i}\alpha_{i}\wedge\beta_{i}\bigr)=0$, which forces $\Psi_{p}\bigl(\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\bigr)=0$.
@@
 By linearity, $\Psi_{p}(\xi)=0$, so $\xi\in\ker\Psi_{p}$.
-For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Equation (5.9) gives
-$A:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad B:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
-Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$.
-Set $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and $V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2})$.
-Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{\dim U}\}$ whose classes are a basis of the quotient $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$).
-By Lemma 5.1, at a generic point $p_{1}\in M_{1}$ the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{\dim U}$ are linearly independent.
-Since $A$ is parallel, $\nabla^{g_{1}}A\equiv 0$; evaluating at such a $p_{1}$ yields
-$\sum_{i>s}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0,$
-and the linear independence forces $c_{i}=0$ for all $i>s$.
-An identical construction applied to $M_{2}$ (choosing a basis of $V_{\parallel}$ and completing it to a basis of $V$ such that the remaining vectors form a basis of the quotient $V/V_{\parallel}$) implies $c_{i}=0$ for all $i>t$.
+For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Then (5.7) holds, i.e.
+$\eta_{1}=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
+Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$. Define
+$U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1}),\qquad V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2}),$
+and set
+$s=\dim U_{\parallel},\qquad t=\dim V_{\parallel}.$
+Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ whose classes form a basis of $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$).
+Now apply Lemma 5.1 to the subspace $U\subset\mathcal{H}^{2}(M_{1})$ with its parallel part $U_{\parallel}$ (which has dimension $s$).
+The lemma guarantees a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
+Since $\eta_{1}$ is parallel, $\nabla^{g_{1}}\eta_{1}\equiv 0$; evaluating at $p_{1}$ we obtain
+$\sum_{i=s+1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0.$
+The linear independence at $p_{1}$ forces $c_{i}=0$ for all $i=s+1,\dots,r_{2,1}$.
+An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}$ such that
+$\sum_{i=t+1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(p_{2})=0,$
+and the linear independence of $\{\nabla^{g_{2}}\beta_{i}(p_{2})\}_{i=t+1}^{r_{2,1}}$ forces $c_{i}=0$ for all $i=t+1,\dots,r_{2,1}$.
 Hence, the only possibly non-zero coefficients are those with $i\leq\min(s,t)$.
@@
 $\ker\Psi_{p}\subseteq\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
-The opposite inclusion is immediate: any element of $\mathcal{V}_{2,1}$ that belongs to $\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})$ is a finite sum of simple tensors of parallel forms; consequently its covariant derivative vanishes, and hence it lies in $\ker\Psi_{p}$.
-Thus we have proved
+The opposite inclusion has already been proved above. Thus
 $\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
@@
 An entirely analogous argument for the (1,2)(1,2)‑component gives
-$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)$ (5.11)
-and
 $\dim(\operatorname{im}\widetilde{\Psi}_{p})=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$ (5.9)
-Hence, on the open‑dense set $\mathcal{U}$ where the harmonic evaluations are independent, the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
+Recall that $s=\dim(U_{\parallel})$ and $t=\dim(V_{\parallel})$.
+By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent.
+Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a non‑empty open set $\mathcal{U}_{2}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
+Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a non‑empty open subset of $M=M_{1}\times M_{2}$.
+For any point $p=(p_{1},p_{2})\in\mathcal{U}$ the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, i.e.
+$\dim\operatorname{im}\Psi_{p}=r_{2,1}-\dim\ker\Psi_{p},\qquad\dim\operatorname{im}\widetilde{\Psi}_{p}=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$
+Hence, at every $p\in\mathcal{U}$, the off‑diagonal part of the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
 $\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
-linearly independent off‑diagonal curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra, whence
+linearly independent curvature operators.
+By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra at $p$.
+By Definition 3.2 of the off‑diagonal holonomy algebra $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ as the Lie subalgebra generated by all off‑diagonal curvature components and their covariant derivatives, we therefore have
 $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
-Consider the one-parameter family of connections
-$\nabla^{(s)}=\nabla^{LC}+\frac{s}{2}T.$
-Then the off-diagonal curvature satisfies $R_{\mathrm{off}}(s)=s(\nabla^{LC}T)_{\mathrm{off}}+s^{2}Q_{\mathrm{off}}$.
-Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\mathrm{off}}$ has dimension at least $r^{\sharp}$.
-The key step is to ensure that the rank of $R_{\mathrm{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\mathrm{off}}$.
-The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$.
-On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$.
-The rank $r^{\sharp}$ is established by the harmonic component $\omega_{h}$ of $T$.
-The non-harmonic part $T-\omega_{h}=d\alpha+\delta\beta$ contributes to $A_{1}$, but since the bound is a lower bound, this contribution cannot reduce the rank below $r^{\sharp}$ on the dense set $U$ and is therefore absorbed by the overall estimate $\mathrm{rank}(A_{1})\geq r^{\sharp}$.
-More formally, a systematic cancellation of the off-diagonal contributions by the non-harmonic part $\eta=d\alpha+\delta\beta$ on an open set would require $\nabla\eta=-\nabla\omega_{h}$ locally.
-By the Unique Continuation Principle, this identity would extend globally, implying $\omega_{h}]=0$, which contradicts our hypothesis of a non-trivial mixed class.
-This ensures that the topological information carried by $\omega_{h}$ is rigidly encoded in the curvature.
-To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in Remark 5.6).
-By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), we can select a sequence of real-analytic product metrics $g_{k}=g_{1,k}\oplus g_{2,k}$ converging to $g$ in the $C^{\infty}$ topology and satisfying
-$\dim\mathcal{P}_{k}(M_{i},g_{k})\leq\dim\mathcal{P}_{k}(M_{i},g)\qquad(k=1,2).$
-This selection is possible because the dimension function $g^{\prime}\mapsto\dim\mathcal{P}_{k}(M_{i},g^{\prime})$ is upper semicontinuous; consequently, the set $\{g^{\prime}\mid\dim\mathcal{P}_{k}(M_{i},g^{\prime})\leq\dim\mathcal{P}_{k}(M_{i},g)\}$ is open in the $C^{\infty}$ topology.
-Real-analytic metrics are dense in this open set by Greene–Jacobowitz.
-For such a sequence we have $\dim\mathcal{K}(g_{k})\leq\dim\mathcal{K}(g)$, and therefore
-$r^{\sharp}(g_{k})\;:=\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}(g_{k})\;\geq\;r^{\sharp}(g).$
-For each fixed $k$, the curvature tensor $R_{\mathrm{off},k}(s)$ of the connection $\nabla^{(s)}=\nabla^{LC}(g_{k})+\frac{s}{2}T$ is real-analytic in $(s,p)$.
-On the dense open set $U$ furnished by Lemma 5.1 we have $\operatorname{rank}(A_{1,k})=r^{\sharp}(g_{k})$, where $A_{1,k}=\frac{1}{2}(\nabla^{LC}(g_{k})T)_{\mathrm{off}}$.
-Hence there exists an $r^{\sharp}(g_{k})\times r^{\sharp}(g_{k})$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ with $M_{k}(s)=s^{r^{\sharp}(g_{k})}D_{k}(s)$ and $D_{k}(0)\neq 0$.
-By the Identity Principle for Analytic Functions, $D_{k}(s)$ does not vanish identically and its zero set is discrete; consequently, for each $k$ there exists an open–dense subset $U_{k}\subseteq U$ such that
-$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k})\qquad\forall p\in U_{k}.$
-Let $W:=\bigcap_{k=1}^{\infty}U_{k}$.
-Since $M$ is compact, it carries a complete Riemannian metric; in a complete metric space, a countable intersection of open dense subsets is dense.
-Thus $W$ is dense in $U$, hence dense in $M$.
-For every $p\in W$ and every $k$ we have
-$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k}).$
-The rank of a continuous family of linear maps is lower semicontinuous with respect to uniform convergence [9, Warner].
-Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, for every $p\in W$ we obtain
-$\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;\liminf_{k\to\infty}\operatorname{rank}R_{\mathrm{off},k}(p)\;\geq\;\liminf_{k\to\infty}r^{\sharp}(g_{k})\;\geq\;r^{\sharp}(g).$
-The set $\{p\in M\mid\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)\}$ is open (by lower semicontinuity) and contains the dense subset $W$; therefore it is all of $M$.
-Thus $\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)$ for every $p\in M$.
-Finally, by Definition 3.2 the off‑diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ is the Lie subalgebra generated by the off‑diagonal curvature operators and their covariant derivatives at $p$.
-In particular, it contains the linear span of the off‑diagonal curvature operators $R_{\mathrm{off}}(p)$.
-Therefore,
-$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;r^{\sharp}(g)\qquad\forall p\in M,$
-which is the desired bound.
+Now let $p_{0}$ be any point in $\mathcal{U}$.
+Because the manifold $M$ is connected, the holonomy algebras at different points are isomorphic via parallel transport along curves.
+In particular, the dimension of the off‑diagonal holonomy algebra is constant on $M$.
+Hence the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$, which holds at $p_{0}$, must hold at every point $p\in M$ and no density requirement is needed.
+This completes the proof of the lower bound.
 The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$.
@@
 Remark 5.4 (Geometric Significance of Invariance).
-The invariance of the lower bound $r^{\sharp}$ is a non‑trivial and geometrically significant property. The bound depends only on two pieces of data:
+The invariance of the lower bound $r^{\sharp}$ is a non‑trivial property that holds under the Assumption 2.2. In those cases, the bound depends only on:
   • the mixed class $\omega]_{\mathrm{mixed}}$,
   • the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics.
-Explicitly,
-$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
-where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (resp.
-$\mathcal{V}_{1,2}$) denotes the linear span of the harmonic representatives of the $(2,1)$- (resp.
-$(1,2)$-) Künneth component of $\omega]$.
-Consequently, $r^{\sharp}$ is invariant under any deformation of the product metric $g=g_{1}\oplus g_{2}$ that preserves the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ ($k=1,2$).
-Such deformations include, for instance, any variation of $g_{i}$ inside the connected component of the moduli space of metrics with a fixed set of parallel forms.
-This stability is especially relevant when at least one factor $M_{i}$ possesses special holonomy (e.g., Calabi–Yau manifolds or flat tori), where the spaces $\mathcal{P}^{k}(M_{i})$ are non‑trivial and the metric $g_{i}$ admits large families of deformations that leave $\mathcal{P}^{k}(M_{i})$ unchanged.
-Thus, $r^{\sharp}$ quantifies the part of the topological “mixing strength” of $\omega]$ that cannot be “absorbed” by metric deformations respecting the parallel‑form strata; it captures a genuinely topological obstruction to reducible holonomy.
+Consequently, $r^{\sharp}$ is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
+This stability is relevant for manifolds with special holonomy where $\mathcal{P}_{k}(M_{i})$ are non‑trivial.
 Remark 5.5 (Forced irreducibility and quantified off‑diagonal holonomy).
@@
 In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
-Remark 5.6 (Analyticity as a tool of proof).
-The use of real‑analytic metrics in the proof of Theorem 5.2 is purely instrumental.
-By the density theorem of Greene–Jacobowitz [4] and the upper semicontinuity of the dimension of the space of parallel forms, one can approximate any smooth product metric $g=g_{1}\oplus g_{2}$ in the $C^{\infty}$-topology by a sequence of real‑analytic product metrics $\{g^{(k)}\}$ satisfying $\dim\mathcal{P}_{\ell}(M_{i},g^{(k)})\leq\dim\mathcal{P}_{\ell}(M_{i},g)$ for $\ell=1,2$.
-For each analytic metric $g^{(k)}$ the curvature tensor $R_{\mathrm{off},k}(s)$ depends analytically on the deformation parameter $s$ and on the point $p\in M$; consequently the rank of the off‑diagonal curvature is lower‑semicontinuous and can be controlled by the behaviour of analytic minors (Identity Principle).
-Passing to the limit $k\to\infty$ via the lower‑semicontinuity of matrix rank with respect to uniform convergence [9] yields the desired bound for the original $C^{\infty}$ metric $g$.
-Thus the analytic approximation does not impose any restriction on the result—it merely provides a convenient technical tool for extending the estimate from a generic dense subset to the whole manifold.
+Remark 5.6 (Analytic approximation as a tool of proof).
+The proofs of Lemma 5.1 and Theorem 5.2 make essential use of real‑analytic approximations of the smooth metric.
+This technique does not restrict the generality of the results, because real‑analytic metrics are dense in the $C^{\infty}$ topology (Greene–Jacobowitz [4]) and the relevant quantities – the dimension of the space of parallel forms, the harmonic representatives of fixed cohomology classes, and the covariant derivatives of those harmonic forms – depend continuously on the metric in the appropriate topologies.
+In Lemma 5.1 the reduction to an analytic metric allows us to apply the identity principle for analytic functions: if an analytic section of a vector bundle vanishes on an open set, it vanishes identically.
+This principle is used to show that a relation among the covariant derivatives of harmonic forms on an open set forces the existence of a parallel combination of the forms, contradicting the hypothesis of linear independence modulo parallel forms.
+For a smooth metric, we approximate by analytic metrics, prove the existence of a non‑empty open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
+In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1.
+Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows first on that open set, and then on the whole manifold because the dimension of the holonomy algebra is constant on connected components (Ambrose–Singer [2]).
+Thus analyticity serves solely as a convenient tool to handle the delicate rank conditions in Lemma 5.1; the final statements are valid for arbitrary smooth product metrics and do not require any analyticity assumption on the metric or the torsion.
 6. Illustrative Examples
@@
 $\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
-7. Significance and Applications to Physics and STEM Disciplines
-Connections with torsion and product structures appear naturally in a wide range of mathematical and physical models.
-In string theory, supergravity, and related frameworks, metric connections with totally skew-symmetric torsion play a central role in flux compactifications, generalized geometries, and effective field theories on product manifolds such as $M_{4}\times X$.
-In these settings, the holonomy of the canonical connection governs the coupling between the external (space-time) and internal (compactified) factors, and irreducibility prevents the decoupling of physical degrees of freedom.
-Our lower bound provides an efficient mechanism for establishing such irreducibility directly from global invariants - particularly the mixed tensor rank and the corrected quantity $r^{\sharp}$ - without computing curvature tensors or higher covariant derivatives.
-This is especially relevant in high-dimensional backgrounds where explicit curvature calculations are notoriously difficult or mathematically intractable.
-A similar phenomenon arises in gauge theory, condensed matter, and materials science, where product manifolds model layered, composite, or multi-component systems.
-The holonomy algebra encodes the geometric interactions between these components; detecting whether such interactions are essential (irreducible) or merely apparent becomes a fundamental structural question.
-The criterion developed in this paper identifies irreducibility based solely on the existence and non-parallelity of mixed differential forms, providing a geometric diagnostic that is robust under perturbations of the underlying metric and torsion.
-Applications also extend to geometric analysis and computational modeling.
-In multi-scale PDE systems defined on product domain (such as transport-diffusion equations, elasticity models, or geometric flows) the holonomy of a natural geometric connection determines how different subsystems interact and whether the governing equations admit a decoupled formulation.
-Our lower bound offers a rigorous obstruction to such decoupling, enabling researchers to understand when full coupling is intrinsic to the geometry.
-This is particularly useful in numerical simulations, where direct computation of curvatures is expensive and unstable, while global invariants such as parallel-form spaces are straightforward to detect algorithmically.
-Overall, the ability to certify irreducible holonomy through topological or geometric invariants provides a powerful tool across physics, geometry, and applied sciences.
-It eliminates the need for explicit curvature computations and transforms a classically analytic problem into a global geometric one, opening the door to applications in any setting where torsion, product structures, and holonomy play a foundational role.
-8. Concluding Remarks and Perspectives
-We have shown that for every compact and orientable product manifold, endowed with any product metric $g$ and for every metric connection satisfying $\nabla^{\mathrm{LC}}T\neq 0$, with 3-form torsion representing a mixed cohomology class via its harmonic projection $\omega_{h}$, the off-diagonal holonomy dimension admits a quantified lower bound
-$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K},$
-where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\tilde{\Psi}_{p}$ is the sum of the kernels of the linear maps $\Psi_{p},\tilde{\Psi}_{p}$ that encode the obstruction to off-diagonal curvature.
-The resulting lower bound is metric-invariant modulo parallel-form preserving deformations, thus providing the first computable obstruction to reducible holonomy on product manifolds.
-This quantifies how the topological “mixing strength” of $\omega]$ forces geometric “entanglement” across the product splitting, modulo the parallel subspace contained within the mixed class.
-The examples above illustrate the computation of $r^{\sharp}$ in simple geometric settings.
-When the torsion is purely harmonic and the parallel-form spaces are minimal, the bound is often saturated in explicit constructions (e.g., on flat tori with parallel torsion).
-It is natural to conjecture that equality holds generically for purely harmonic torsion, while strict inequality may occur when the coexact or exact parts of $T$ generate additional off-diagonal curvature directions, or when the parallel-form spaces are larger.
-A systematic classification of the cases where the bound is sharp remains an interesting direction for future work.
-A further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].
+7. Conclusion
+We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$—specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings with rank‑one mixed classes—and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound
+$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=r-\dim\mathcal{K},$
+where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\operatorname{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the sum of the kernels of the linear obstruction maps.
+The bound is metric‑invariant under deformations preserving the parallel‑form strata $\mathcal{P}_{k}(M_{i})$.
+This provides the first computable obstruction to reducible holonomy on product manifolds within the admissible class.
+The examples in Section 6 illustrate both non‑trivial and trivial cases, while Remark 2.4 shows by explicit construction that the bound fails to be well‑defined without the admissibility hypothesis, establishing sharpness of our assumptions.
+When the torsion is purely harmonic and the parallel‑form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion).
+It is natural to conjecture that equality holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off‑diagonal curvature or when the parallel‑form spaces are larger.
+A systematic classification of sharp cases remains an interesting open problem.
+A further direction is to explore extensions beyond the admissible setting.
+Imposing Kruskal‑rank constraints on harmonic representatives could yield intrinsic mixed factor spaces even when $b_{2}(M_{i})>1$, potentially treating Calabi–Yau threefolds or $G_{2}\times S^{1}$ backgrounds.
+Such extensions would require new algebraic tools and are left for future investigation.
+Another further interesting direction for future work could be to explore the implications of these results in the context of deformations of complex structures, a field pioneered by the works of Kodaira [6] and Kodaira-Spencer [7].
 Compactness and Hodge theory.

### v11

@@
 Abstract.
-We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting and the mixed class has tensor rank $1$.
-For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove
+We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting.
+For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove that on a non-empty open subset $\mathcal{U}\subset M$,
 $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
 with $\mathcal{K}$ an intrinsically defined obstruction space.
-The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides a sharp obstruction to reducible holonomy.
+The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides an obstruction to reducible holonomy.
 Counterexamples show the hypothesis is optimal.
@@
 1. Introduction
-Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$.
+Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold [^1] As is customary in Riemannian geometry, all manifolds are assumed to be connected unless otherwise stated.
+endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$.
 We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
@@
   • (B)
-    $\dim M=4$ with a $(3,1)$ splitting and the mixed class has total tensor rank $1$.
-For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies
+    $\dim M=4$ with a $(3,1)$ splitting.
+For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies, on a non-empty open subset $\mathcal{U}\subset M$,
 $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
@@
     $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
+[^2] This is not an additional condition, but holds automatically: $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\geq 1$ because $\omega]_{\mathrm{mixed}}\neq 0$ (not trivial mixed class), while $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\leq 1$ since $S^{1}$ is the unique compact connected oriented $1$-manifold, with $\dim H^{1}(S^{1})=1$ and $H^{2}(S^{1})=0$; hence equality holds.
 Lemma 2.3 (Uniqueness of Mixed Subspaces).
@@
 is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$.
-Case (B): $(3,1)$ or $(1,3)$ with $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
+Case (B): $(3,1)$ or $(1,3)$.
 Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric).
@@
 Uniqueness follows as in Case (A) because $\gamma$ is non–zero.
-Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ depends only on $\omega]$.
+Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ is intrinsically defined.
 The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined.
@@
 and we set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
+Remark 2.6 (Anticipatory notation).
+The notation $\ker\Psi_{p}$” is anticipatory: in the proof of Theorem 5.2, we construct obstruction maps $\Psi_{p}:V_{2,1}\to\mathfrak{so}(T_{p}M)$ and $\widetilde{\Psi}_{p}:V_{1,2}\to\mathfrak{so}(T_{p}M)$ via covariant derivatives of the torsion form.
+A key step in the proof is to verify that the kernels of these maps coincide with the intersections defined above.
+The induced quotient maps $V_{2,1}/\ker\Psi_{p}\to\mathfrak{so}(T_{p}M)$ are then shown to be injective on a non-empty open subset, establishing the lower bound.
 3. Off-Diagonal Curvature and Holonomy Complexity
@@
 $R_{\mathrm{off}}(X,Y)Z:=P_{1}R(X,Y)(P_{2}Z)+P_{2}R(X,Y)(P_{1}Z).$ (3.1)
-Definition 3.2 (Off-diagonal holonomy algebra).
+Definition 3.2 (Off-diagonal holonomy component).
 Let $(M,g)$ be a Riemannian product manifold with orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$, and let $\nabla^{C}$ be a metric connection on $TM$ with totally skew-symmetric torsion.
-Let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
-$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z),$
-where $P_{i}:T_{p}M\to V_{i}$ are the orthogonal projections.
-Let $\mathcal{G}_{p}:=\{\nabla^{k}R^{C}|_{p}(X_{1},\dots,X_{k};Y,Z)\mid k\geq 0,X_{i},Y,Z\in T_{p}M\}$ be the set of Ambrose–Singer curvature generators of $\mathfrak{hol}_{p}(\nabla^{C})$.
-Define the off-diagonal curvature span
-$\mathcal{S}_{p}^{\mathrm{off}}:=\operatorname{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(A)\mid A\in\mathcal{G}_{p}\}\subset\mathfrak{so}(T_{p}M).$
-The off-diagonal holonomy algebra is the Lie subalgebra obtained by intersecting $\mathfrak{hol}_{p}(\nabla^{C})$ with the Lie algebra generated by $\mathcal{S}_{p}^{\mathrm{off}}$ in $\mathfrak{so}(T_{p}M)$:
-$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\bigl\langle\mathcal{S}_{p}^{\mathrm{off}}\bigr\rangle_{\mathrm{Lie}}\cap\mathfrak{hol}_{p}(\nabla^{C})\subset\mathfrak{hol}_{p}(\nabla^{C}).$
-Here $\langle\cdot\rangle_{\mathrm{Lie}}$ denotes the Lie algebra generated in $\mathfrak{so}(T_{p}M)$, and the intersection ensures that $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is canonically a subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$.
-Equivalently, it is the smallest Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ whose off-diagonal components contain $\mathcal{S}_{p}^{\mathrm{off}}$.
+Let $P_{i}:T_{p}M\to V_{i}$ be the orthogonal projections and let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
+$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z).$
+Define the space of off-diagonal skew-symmetric endomorphisms as the image of this projection:
+$\mathfrak{so}^{\mathrm{off}}(T_{p}M):=\operatorname{Im}(\Pi_{\mathrm{off}})=\{A\in\mathfrak{so}(T_{p}M)\mid A(V_{1})\subset V_{2},\,A(V_{2})\subset V_{1}\}.$
+The off-diagonal holonomy component is the intersection of the holonomy algebra with this space:
+$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M).$
+Equivalently, $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{A\in\mathfrak{hol}_{p}(\nabla^{C})\mid\Pi_{\mathrm{off}}(A)=A\}$ consists of those holonomy elements that are purely off-diagonal.
+Remark 3.3.
+Although $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is a vector subspace of $\mathfrak{hol}_{p}(\nabla^{C})$, it is not closed under the Lie bracket, since $\mathfrak{so}^{\mathrm{off}},\mathfrak{so}^{\mathrm{off}}]\subset\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$.
+Nevertheless, $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})>0$ implies that the holonomy representation is irreducible with respect to the splitting $V_{1}\oplus V_{2}$.
 By the Ambrose–Singer theorem [2], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor.
@@
 Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$.
-Then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
+If $r-s\leq 1$ (condition satisfied by admissible product manifolds: $U$ is the span of a minimal harmonic decomposition, so $\dim U=r_{2,1}\leq 1$, hence $r-s\leq 1$; see Definition 4.1, Definition 2.2 and Lemma 2.3), then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
 $\bigl(\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}\bigr)$
@@
 2.
-The real‑analytic case.
+Admissible manifolds.
 Assume now that $g_{1}$ is real‑analytic.
@@
 $\sum_{i=1}^{r-s}c_{i}(p)\bigl(\nabla^{g_{1}}\alpha_{s+i}\bigr)(p)=0\qquad\forall p\in V.$ (5.1)
-Define the analytic $2$-form $\eta:=\sum_{i=1}^{r-s}c_{i}\alpha_{s+i}$ on $V$.
-From (5.1) we have $\nabla^{g_{1}}\eta=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}+\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}$.
-Let $V\subset M_{1}$ be the non–empty open subset fixed above.
-Since the mixed cohomology class on $M_{1}\times M_{2}$ is assumed to be non–trivial, its harmonic representative has a non–vanishing component in $H^{2}(M_{1})\otimes H^{1}(M_{2})$.
-In particular, the projection onto $H^{2}(M_{1})$ is non–zero.
-Hence the harmonic forms $\alpha_{s+1},\dots,\alpha_{r}$ cannot all vanish identically on $M_{1}$.
-Let $W\subset M_{1}$ denote the complement of their common zero set.
-By real–analyticity, $W$ is a non–empty open subset of $M_{1}$.
-Restricting to the non‑empty open set $V^{\prime}:=V\cap W$, we have on $V^{\prime}$
-$\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}=0.$
-Because the $\alpha_{s+i}(p)$ are linearly independent at every $p\in V^{\prime}$, this forces $dc_{i}=0$ on $V^{\prime}$ for each $i$.
-Thus the functions $c_{i}$ are constant on each connected component of $V^{\prime}$.
-By analyticity, they extend to constants on the whole connected open set $V$ (if two analytic functions coincide on an open set, they coincide everywhere on the connected component).
-Consequently, the $c_{i}$ are constant on $V$.
-Thus there exist constants $c_{1},\dots,c_{r-s}\in\mathbb{R}$, not all zero, such that
-$\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=0\qquad\text{on }V.$
-Define the global harmonic $2$-form
-$\zeta:=\sum_{i=1}^{r-s}\bar{c}_{i}\alpha_{s+i}\in\mathcal{H}^{2}(M_{1}).$
-Then $\nabla^{g_{1}}\zeta=0$ on $V$; i.e., $\zeta$ is parallel on the open set $V$.
-Because both $\zeta$ and the condition $\nabla^{g_{1}}\zeta=0$ are real-analytic, the unique-continuation principle [1] implies that $\nabla^{g_{1}}\zeta=0$ on the whole connected component of $M_{1}$ containing $V$, hence on all of $M_{1}$.
-Therefore $\zeta$ is a globally parallel $2$-form.
-But this contradicts the hypothesis that $\alpha_{s+1},\dots,\alpha_{r}$ are linearly independent modulo $\mathcal{P}_{2}(M_{1},g_{1})$.
-Indeed, $\zeta$ is a non-trivial linear combination of these forms that belongs to $\mathcal{P}_{2}(M_{1},g_{1})$.
-Hence our assumption that $\mathcal{U}=\emptyset$ is impossible, and we conclude that $\mathcal{U}$ is a non-empty open subset of $M_{1}$.
+Since we have to consider admissible manifolds (Definition 2.2 and Lemma 2.3) we have two cases:
+Case $r=s$: There are no non‑parallel forms to consider; the lemma holds vacuously.
+Case $r=s+1$: There exists a unique non‑parallel form $\alpha$ and a single analytic function $c$ such that (5.1) reads $c(p)(\nabla^{g_{1}}\alpha)_{p}=0$ for all $p\in V$.
+If $c(p_{0})\neq 0$ for some $p_{0}\in V$, then by analyticity $c(p)\neq 0$ on an open subset $V^{\prime}\subset V$, whence $(\nabla^{g_{1}}\alpha)_{p}=0$ on $V^{\prime}$.
+By the unique‑continuation principle for analytic sections (Aronszajn [1]; see also [8]), $\nabla^{g_{1}}\alpha\equiv 0$ on the whole connected component containing $V^{\prime}$, and hence on all of $M_{1}$.
+This contradicts the assumption that $\alpha$ is non‑parallel modulo $\mathcal{P}_{2}(M_{1},g_{1})$.
+Therefore $c\equiv 0$ on $V$, contradicting the choice of the $c_{i}$’s.
+Hence $\mathcal{U}\neq\emptyset$.
 3.
@@
 ∎
+The proof uses only properties of harmonic forms that hold for any degree $k$: analytic approximation, semicontinuity of the dimension of parallel forms, $C^{1}$ convergence, and unique continuation.
+Therefore, the same result holds for $\mathcal{H}^{1}(M_{1},g_{1})$, as well as for $\mathcal{H}^{1}(M_{2},g_{2})$ and $\mathcal{H}^{2}(M_{2},g_{2})$.
 Theorem 5.2 (Topological Lower Bound for Off-Diagonal Curvature).
@@
 and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
-Then, for every $p\in M$,
+Then, there exists a non-empty open subset $\mathcal{U}\subset M$ such that for every $p\in\mathcal{U}$,
 $\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
@@
 Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
-Consider the curvature formula (2.5).
+Consider the curvature formula (2.5):
+$R^{C}=R^{LC}+\frac{1}{2}(\nabla^{LC}T)+\frac{1}{4}Q_{T}.$
 The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts.
-Our aim is to show that the off‑diagonal holonomy algebra contains at least $r^{\sharp}$ linearly independent directions.
-These directions will be produced by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the harmonic part $\omega_{h}$ alone.
-We must verify that the remaining contributions—the part of the linear term coming from $\eta$ and the quadratic term $Q_{T}$—cannot interfere with this lower bound.
-Suppose that on some non‑empty open set $V\subset M$, the derivative $\nabla^{LC}\eta$ cancelled enough off‑diagonal directions from $\nabla^{LC}\omega_{h}$ to reduce the rank of $(\nabla^{LC}T)_{\mathrm{off}}$ below $r^{\sharp}$.
-This would impose linear relations among the covariant derivatives of the harmonic components $\alpha_{i}\wedge\beta_{i}$ on $V$.
-Applying the analytic‑approximation technique employed in Lemma 5.1 (see also Remark 5.6), we may work with real‑analytic metrics and forms.
-The relation then becomes an equality between real‑analytic sections, so by the unique continuation principle for analytic sections (Aronszajn [1]) it extends to the whole connected component containing $V$, hence globally.
-Since the tensors $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})$ are pointwise linearly independent on the open set $\mathcal{U}$ provided by Lemma 5.1, evaluating the global relation at any point of $\mathcal{U}$ forces all coefficients to vanish, contradicting the existence of a non‑trivial cancellation.
-Hence, on a (possibly smaller) non‑empty open set, the linear term $(\nabla^{LC}T)_{\mathrm{off}}$ still exhibits at least $r^{\sharp}$ independent off‑diagonal directions originating from $\omega_{h}$.
-Regarding the quadratic term $Q_{T}$, observe that the $r^{\sharp}$ independent operators produced by the linear term belong to the off‑diagonal holonomy algebra by the Ambrose–Singer theorem [2].
-The quadratic term contributes additional operators to the same algebra.
-Since the dimension of a vector space generated by a set of vectors is at least the rank of any linearly independent subset, adding further generators cannot reduce the dimension below $r^{\sharp}$.
-Consequently, the lower bound obtained from the linear term persists for the full curvature $R^{C}$.
-Thus, to establish $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ it suffices to analyse the linear term generated by the harmonic component $\omega_{h}$.
+Taking off-diagonal components:
+$R^{C}_{\mathrm{off}}=R^{LC}_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\eta)_{\mathrm{off}}+\frac{1}{4}(Q_{T})_{\mathrm{off}}.$
+Our aim is to show that $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}:=r-\dim\mathcal{K}$ on a non-empty open subset of $M$.
+We first establish that to prove this bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic term $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ alone.
+Suppose that at some point $p\in M$ we have constructed operators $A_{1},\ldots,A_{r^{\sharp}}\in\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ that are linearly independent and obtained from evaluations of $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$.
+By the curvature formula, each such operator is a value of $R^{C}_{\mathrm{off}}$, since $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ appears as one of the terms in $R^{C}_{\mathrm{off}}$.
+Let
+$S_{p}^{\mathrm{off}}:=\mathrm{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(R^{C}(X,Y)):X,Y\in T_{p}M\}$
+be the span of all off-diagonal curvature values at $p$.
+By the curvature formula (2.5), the covariant derivative $(\nabla^{LC}_{X}T)(Y,Z,\cdot)$ appears in the operator $R^{C}(X,Y)$ after raising indices; hence the operators $A_{1},\ldots,A_{r^{\sharp}}$ constructed from evaluations of $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ are indeed components of curvature values, and therefore belong to $S_{p}^{\mathrm{off}}$.
+Since they are linearly independent, we have $\dim S_{p}^{\mathrm{off}}\geq r^{\sharp}$.
+The terms $R^{LC}_{\mathrm{off}}$, $(\nabla^{LC}\eta)_{\mathrm{off}}$, and $(Q_{T})_{\mathrm{off}}$ contribute additional values to $S_{p}^{\mathrm{off}}$, but cannot reduce its dimension below $r^{\sharp}$.
+By the Ambrose–Singer theorem [2], $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ contains $S_{p}^{\mathrm{off}}$ as the off-diagonal part of the algebra generated by curvature values, whence $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$.
+Thus, to establish the desired lower bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic component $\omega_{h}$.
 To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$.
@@
 $\eta_{1}:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.7)
+(Note that the coefficients $\{c_{i}\}$ are the same in both conditions, being the components of the fixed element $\xi=\sum c_{i}\alpha_{i}\otimes\beta_{i}$ in the tensor product decomposition.)
 Conversely, if (5.7) holds, then $\nabla^{LC}\bigl(\sum_{i}c_{i}\alpha_{i}\wedge\beta_{i}\bigr)=0$, which forces $\Psi_{p}\bigl(\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\bigr)=0$.
@@
 linearly independent curvature operators.
-By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra at $p$.
-By Definition 3.2 of the off‑diagonal holonomy algebra $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ as the Lie subalgebra generated by all off‑diagonal curvature components and their covariant derivatives, we therefore have
+By the Ambrose–Singer theorem [2], the holonomy algebra $\mathfrak{hol}_{p}(\nabla^{C})$ is generated by curvature values.
+The off-diagonal component $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ (Definition 3.2) contains all off-diagonal projections of elements of $\mathfrak{hol}_{p}(\nabla^{C})$.
+Since the $r^{\sharp}$ operators constructed above are off-diagonal curvature operators, they belong to $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$, and therefore
 $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
-Now let $p_{0}$ be any point in $\mathcal{U}$.
-Because the manifold $M$ is connected, the holonomy algebras at different points are isomorphic via parallel transport along curves.
-In particular, the dimension of the off‑diagonal holonomy algebra is constant on $M$.
-Hence the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$, which holds at $p_{0}$, must hold at every point $p\in M$ and no density requirement is needed.
-This completes the proof of the lower bound.
-The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$.
+In particular, if $r^{\sharp}>0$, the holonomy representation is irreducible with respect to $V_{1}\oplus V_{2}$.
+The lower bound depends only on $\omega]_{\mathrm{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; it is therefore invariant under any product-metric deformation that preserves these strata.
 ∎
@@
 In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1.
-Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows first on that open set, and then on the whole manifold because the dimension of the holonomy algebra is constant on connected components (Ambrose–Singer [2]).
+Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows on that open set.
 Thus analyticity serves solely as a convenient tool to handle the delicate rank conditions in Lemma 5.1; the final statements are valid for arbitrary smooth product metrics and do not require any analyticity assumption on the metric or the torsion.
@@
 7. Conclusion
-We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$—specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings with rank‑one mixed classes—and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound
+We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$, specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings, and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound on a non-empty open subset $\mathcal{U}\in M$
 $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=r-\dim\mathcal{K},$

### v12

@@
 Abstract.
-We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting.
-For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove that on a non-empty open subset $\mathcal{U}\subset M$,
-$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
+We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on product manifolds where both factors are compact oriented surfaces.
+For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove that on a non-empty open subset $\mathcal{V}\subset M$,
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq\min\bigl(\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},\;1\bigr),$
 with $\mathcal{K}$ an intrinsically defined obstruction space.
@@
 1. Introduction
-Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold [^1] As is customary in Riemannian geometry, all manifolds are assumed to be connected unless otherwise stated.
-endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$.
+Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$.
 We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
 A subtle but crucial point is that the subspaces spanned by a minimal harmonic decomposition are uniquely determined only under topological restrictions on the factors.
-Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e.:
-  • (A)
-    both factors are compact oriented surfaces;  or
-  • (B)
-    $\dim M=4$ with a $(3,1)$ splitting.
-For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies, on a non-empty open subset $\mathcal{U}\subset M$,
-$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
+Consequently, our main result holds for admissible product manifolds (Definition 2.2), i.e., where both factors are compact oriented surfaces.
+For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies, on a non-empty open subset $\mathcal{V}\subset M$,
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\min\bigl(\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},\;1\bigr),$
 where the obstruction space $\mathcal{K}$ is intrinsically defined via intersections of the mixed factor spaces $\mathcal{V}_{2,1}$, $\mathcal{V}_{1,2}$ with the parallel‑form strata $\mathcal{P}_{k}(M_{i})$.
@@
 2.1. Intrinsic Definition of Mixed Cohomology Subspaces
-We now address a subtle but crucial point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general.
+We now address a subtle but fundamental point: the subspaces spanned by the factors in a minimal harmonic decomposition are not uniquely determined in general.
 To obtain a well–defined geometric invariant we must impose conditions on the topology of the factors.
 Definition 2.2 (Admissible Product Manifolds).
-A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies one of the following:
-  • (A)
-    both factors $M_{1},M_{2}$ are compact oriented surfaces;
-  • (B)
-    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
-[^2] This is not an additional condition, but holds automatically: $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\geq 1$ because $\omega]_{\mathrm{mixed}}\neq 0$ (not trivial mixed class), while $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\leq 1$ since $S^{1}$ is the unique compact connected oriented $1$-manifold, with $\dim H^{1}(S^{1})=1$ and $H^{2}(S^{1})=0$; hence equality holds.
+A compact oriented product manifold $M=M_{1}\times M_{2}$ is called admissible if it satisfies the following:
+  • both factors $M_{1}$, $M_{2}$ are compact oriented surfaces.
 Lemma 2.3 (Uniqueness of Mixed Subspaces).
@@
 Proof.
-We treat the two admissible cases separately.
-Case (A): $\dim M_{i}=2$. For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
+For a compact oriented surface $F^{2}$ we have $b_{2}(F)=1$ and $b_{1}(F)=2g(F)$. Let $\operatorname{vol}_{F}$ be the unique (up to scale) $g_{F}$–harmonic 22–form. Then
 $\mathcal{H}^{2}(F)=\mathbb{R}\cdot\operatorname{vol}_{F}.$
@@
 $\mathcal{V}_{2,1}=\mathcal{H}^{2}(M_{1})\otimes\operatorname{span}\{\beta\}$
-is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$.
-Case (B): $(3,1)$ or $(1,3)$.
-Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric).
-Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$.
-Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$.
-Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
-$T=\alpha\otimes\gamma,\qquad\alpha\in\mathcal{H}^{2}(M_{1}).$
-Uniqueness follows as in Case (A) because $\gamma$ is non–zero.
-Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ is intrinsically defined.
-The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined.
-∎
+is intrinsically defined. The same argument applied symmetrically yields the intrinsic definition of $\mathcal{V}_{1,2}$. ∎
 Remark 2.4 (Failure of invariance in general).
@@
 A key step in the proof is to verify that the kernels of these maps coincide with the intersections defined above.
-The induced quotient maps $V_{2,1}/\ker\Psi_{p}\to\mathfrak{so}(T_{p}M)$ are then shown to be injective on a non-empty open subset, establishing the lower bound.
+The induced quotient maps $V_{2,1}/\ker\Psi_{p}\to\mathfrak{so}(T_{p}M)$ are then shown to be injective on a dense open subset, establishing the lower bound.
 3. Off-Diagonal Curvature and Holonomy Complexity
@@
 Although $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is a vector subspace of $\mathfrak{hol}_{p}(\nabla^{C})$, it is not closed under the Lie bracket, since $\mathfrak{so}^{\mathrm{off}},\mathfrak{so}^{\mathrm{off}}]\subset\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$.
-Nevertheless, $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})>0$ implies that the holonomy representation is irreducible with respect to the splitting $V_{1}\oplus V_{2}$.
+Nevertheless, $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})>0$ implies that the holonomy representation is irreducible with respect to the splitting $V_{1}\oplus V_{2}$, therefore, we have global irreducibility.
 By the Ambrose–Singer theorem [3], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor.
@@
 Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$.
-If $r-s\leq 1$ (condition satisfied by admissible product manifolds: $U$ is the span of a minimal harmonic decomposition, so $\dim U=r_{2,1}\leq 1$, hence $r-s\leq 1$; see Definition 4.1, Definition 2.2 and Lemma 2.3), then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
+If $r-s\leq 1$ (condition satisfied by admissible product manifolds: $U$ is the span of a minimal harmonic decomposition, so $\dim U=r_{2,1}\leq 1$, hence $r-s\leq 1$; see Definition 4.1, Definition 2.2 and Lemma 2.3), then there exists a dense open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
 $\bigl(\nabla^{g_{1}}\alpha_{s+1},\dots,\nabla^{g_{1}}\alpha_{r}\bigr)$
@@
 $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}\longrightarrow\nabla^{g_{1}}\alpha_{i}\qquad(i=s+1,\dots,r).$
-If we prove that for each $k$ there exists a non‑empty open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a non‑empty open set for the original metric $g_{1}$.
+If we prove that for each $k$ there exists a dense open set $\mathcal{U}^{(k)}\subset M_{1}$ where the $\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)}$ are linearly independent, then the $C^{1}$ convergence guarantees that, for large enough $k$, the same linear independence holds on a dense open set for the original metric $g_{1}$.
 Indeed, linear independence is equivalent to the non‑vanishing of at least one $(r-s)\times(r-s)$ minor of the matrix of components; each minor depends continuously on the metric and the forms, so a minor that is non‑zero on an open set for $g_{1}^{(k)}$ remains non‑zero on an open set for $g_{1}$ when $k$ is sufficiently large.
@@
 If $c(p_{0})\neq 0$ for some $p_{0}\in V$, then by analyticity $c(p)\neq 0$ on an open subset $V^{\prime}\subset V$, whence $(\nabla^{g_{1}}\alpha)_{p}=0$ on $V^{\prime}$.
-By the unique‑continuation principle for analytic sections (Aronszajn [1]; see also [8]), $\nabla^{g_{1}}\alpha\equiv 0$ on the whole connected component containing $V^{\prime}$, and hence on all of $M_{1}$.
+By the unique‑continuation principle for analytic sections (Aronszajn [1]; for differential forms see [2]), $\nabla^{g_{1}}\alpha\equiv 0$ on the whole connected component containing $V^{\prime}$, and hence on all of $M_{1}$.
 This contradicts the assumption that $\alpha$ is non‑parallel modulo $\mathcal{P}_{2}(M_{1},g_{1})$.
@@
 Hence $\mathcal{U}\neq\emptyset$.
+Moreover, the complement $M_{1}\setminus\mathcal{U}=\{p\in M_{1}:(\nabla^{g_{1}}\alpha)(p)=0\}$ has empty interior: if it contained an open set $V^{\prime}$, then by unique continuation $\nabla^{g_{1}}\alpha\equiv 0$ on all of $M_{1}$, contradicting the hypothesis that $\alpha$ is non-parallel.
+Therefore $\mathcal{U}$ is dense in $M_{1}$.
 3.
 Returning to the original smooth metric.
-As explained in Step 1, the existence of a non‑empty open set of linear independence for a sequence of real‑analytic approximations $g_{1}^{(k)}$ together with the $C^{1}$ convergence of the covariant derivatives guarantees the existence of a non‑empty open set with the same property for the original smooth metric $g_{1}$.
-This completes the proof.
+By Step 2, for each analytic approximation $g_{1}^{(k)}$, the set $\mathcal{U}^{(k)}=\{p:\nabla^{g_{1}^{(k)}}\alpha^{(k)}(p)\neq 0\}$ is dense in $M_{1}$.
+The same conclusion holds for the smooth metric $g_{1}$ by applying the unique continuation principle for harmonic forms on smooth Riemannian manifolds [2] directly: if $M_{1}\setminus\mathcal{U}$ contained an open set $V$, lifting to the universal cover $\tilde{M}_{1}$ yields a globally parallel form $\tilde{\beta}$ with $\tilde{\alpha}|_{\tilde{V}}=\tilde{\beta}|_{\tilde{V}}$.
+Since $\tilde{\alpha}-\tilde{\beta}$ is harmonic and vanishes on $\tilde{V}$, unique continuation [2] implies $\tilde{\alpha}=\tilde{\beta}$, so $\alpha$ is parallel on $M_{1}$, contradicting $\alpha\notin U_{\parallel}$.
+Thus $\mathcal{U}$ is dense.
 ∎
@@
 Let $M=M_{1}\times M_{2}$ be an admissible compact oriented Riemannian product manifold (Definition 2.2) with product metric $g=g_{1}\oplus g_{2}$.
-Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class.
+Let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]$ is a non‑trivial mixed class (Definition 2.1).
 Define the mixed factor subspaces
@@
 and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$.
-Then, there exists a non-empty open subset $\mathcal{U}\subset M$ such that for every $p\in\mathcal{U}$,
-$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
+Then, there exists a non-empty open subset $\mathcal{V}\subset M$ such that for every $p\in\mathcal{V}$,
+$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\;\geq\;r^{\sharp}:=\min\bigl(\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},\;1\bigr).$
 The bound depends only on $\omega]_{\text{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
@@
 The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
-Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
-Consider the curvature formula (2.5):
-$R^{C}=R^{LC}+\frac{1}{2}(\nabla^{LC}T)+\frac{1}{4}Q_{T}.$
-The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts.
-Taking off-diagonal components:
-$R^{C}_{\mathrm{off}}=R^{LC}_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\eta)_{\mathrm{off}}+\frac{1}{4}(Q_{T})_{\mathrm{off}}.$
-Our aim is to show that $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}:=r-\dim\mathcal{K}$ on a non-empty open subset of $M$.
-We first establish that to prove this bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic term $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ alone.
-Suppose that at some point $p\in M$ we have constructed operators $A_{1},\ldots,A_{r^{\sharp}}\in\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ that are linearly independent and obtained from evaluations of $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$.
-By the curvature formula, each such operator is a value of $R^{C}_{\mathrm{off}}$, since $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ appears as one of the terms in $R^{C}_{\mathrm{off}}$.
-Let
-$S_{p}^{\mathrm{off}}:=\mathrm{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(R^{C}(X,Y)):X,Y\in T_{p}M\}$
-be the span of all off-diagonal curvature values at $p$.
-By the curvature formula (2.5), the covariant derivative $(\nabla^{LC}_{X}T)(Y,Z,\cdot)$ appears in the operator $R^{C}(X,Y)$ after raising indices; hence the operators $A_{1},\ldots,A_{r^{\sharp}}$ constructed from evaluations of $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ are indeed components of curvature values, and therefore belong to $S_{p}^{\mathrm{off}}$.
-Since they are linearly independent, we have $\dim S_{p}^{\mathrm{off}}\geq r^{\sharp}$.
-The terms $R^{LC}_{\mathrm{off}}$, $(\nabla^{LC}\eta)_{\mathrm{off}}$, and $(Q_{T})_{\mathrm{off}}$ contribute additional values to $S_{p}^{\mathrm{off}}$, but cannot reduce its dimension below $r^{\sharp}$.
-By the Ambrose–Singer theorem [2], $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ contains $S_{p}^{\mathrm{off}}$ as the off-diagonal part of the algebra generated by curvature values, whence $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$.
-Thus, to establish the desired lower bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic component $\omega_{h}$.
+Set $r^{\sharp}:=\min(r-\dim\mathcal{K},\,1)$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
+Consider the curvature formula (2.5) for $X\in V_{1}$ and $Y\in V_{2}$:
+$R^{C}(X,Y)\;=\;R^{LC}(X,Y)\;+\;\tfrac{1}{2}\bigl[(\nabla^{LC}_{X}T)(Y,\cdot\,,\cdot)^{\sharp}-(\nabla^{LC}_{Y}T)(X,\cdot\,,\cdot)^{\sharp}\bigr]\;+\;\tfrac{1}{4}\,Q_{T}(X,Y).$
+On a Riemannian product the mixed Riemann curvature vanishes: $R^{LC}(X,Y)=0$ whenever $X\in V_{1}$ and $Y\in V_{2}$ (since $\nabla^{LC}_{X}Y=0=[X,Y]$ for vector fields tangent to different factors).
+For the remaining three terms we exploit the rigid pointwise algebraic structure that the surface hypothesis imposes on mixed $3$-forms.
+Since both factors are compact oriented surfaces, $\dim V_{i}=2$, and the exterior power $\Lambda^{2}V_{1}^{*}$ is one-dimensional, spanned by the area form $\operatorname{vol}_{V_{1}}$.
+Every $3$-form of bidegree $(2,1)$ on $M$ can therefore be written at each point $p$ as
+$T^{2,1}_{p}\;=\;h(p)\;\operatorname{vol}_{V_{1}}\wedge\tau_{p},$ (5.4)
+where $h\colon M\to\mathbb{R}$ is a smooth function and $\tau$ is a smooth section of $V_{2}^{*}$ (depending on the full point $p=(p_{1},p_{2})$).
+Likewise, every $(1,2)$-form satisfies $T^{1,2}_{p}=\widetilde{h}(p)\;\widetilde{\sigma}_{p}\wedge\operatorname{vol}_{V_{2}}$ with $\widetilde{\sigma}\in\Gamma(V_{1}^{*})$.
+These representations hold for the full torsion $T$, not merely for its harmonic projection $\omega_{h}$.
+We denote by $J_{i}\colon V_{i}\to V_{i}$ the volume endomorphism of the oriented surface $M_{i}$, characterised by $g(J_{i}u,v)=\operatorname{vol}_{V_{i}}(u,v)$; it is a skew-symmetric isometry with $J_{i}^{2}=-\operatorname{Id}$, and in particular $J_{i}X\perp X$ for every nonzero $X\in V_{i}$.
+Let $\pi_{0}\colon\mathfrak{so}(T_{p}M)\to\mathfrak{g}_{0}:=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$ and $\pi_{1}\colon\mathfrak{so}(T_{p}M)\to\mathfrak{g}_{1}:=\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ denote the projections induced by the $\mathbb{Z}_{2}$-grading of $\mathfrak{so}(T_{p}M)$.
+[…]
+$\pi_{0}\bigl(R^{C}(X,Y)\bigr)\;=\;\lambda(X,Y)\,J_{1}\;+\;\mu(X,Y)\,J_{2}$ (5.5)
+[…]
+$\rho_{Y}\;:=\;(Yh)\,\tau\;+\;h\,\nabla^{g_{2}}_{Y}\!\tau\;=\;\nabla^{g_{2}}_{Y}(h\tau),$ (5.6)
+the restriction to $\operatorname{Hom}(V_{1},V_{2})$ reads
+$\tfrac{1}{2}(\nabla^{LC}_{Y}T^{2,1})(X,Z,\cdot)^{\sharp}\Big|_{Z\in V_{1}}\;=\;-\tfrac{1}{2}\,g(J_{1}X,Z)\;\rho_{Y}^{\sharp}\;\in V_{2}.$ (5.7)
+The third, $\frac{1}{4}Q_{T}(X,Y)$, is also purely off-diagonal (by a bidegree computation), with restriction
+$\tfrac{1}{4}\,Q_{T}(X,Y)\big|_{Z\in V_{1}}\;=\;\tfrac{h^{2}}{4}\;\tau(Y)\;g(X,Z)\;\tau^{\sharp}\;\in V_{2}.$ (5.8)
+To derive 5.8, observe that the contortion tensor $K_{X}Y=\tfrac{1}{2}T(X,Y,\cdot)^{\sharp}$ acts on $Z\in V_{1}$ via the $(2,1)$-component as
+$K_{Y}^{(2,1)}Z=\frac{h}{2}\,\tau(Y)\,J_{1}Z,$
+since $(\operatorname{vol}_{V_{1}}\wedge\tau)(Y,Z,W)=\tau(Y)\,\operatorname{vol}_{V_{1}}(Z,W)$ for $Y\in V_{2}$ and $Z,W\in V_{1}$. The iterated contortion $K_{X}(K_{Y}Z)$ with $X\in V_{1}$ then gives
+$K_{X}(K_{Y}Z)=\frac{h^{2}}{4}\,\tau(Y)\,g(X,Z)\,\tau^{\sharp},$
+where we used $T^{(2,1)}(X,J_{1}Z,W)=h\,\operatorname{vol}_{V_{1}}(X,J_{1}Z)\,\tau(W)=h\,g(J_{1}X,J_{1}Z)\,\tau(W)=h\,g(X,Z)\,\tau(W)$, the last equality following from $J_{1}$ being an isometry.
+Conversely, $K_{Y}(K_{X}Z)=0$: since $K_{X}Z=\tfrac{h}{2}\operatorname{vol}_{V_{1}}(X,Z)\,\tau^{\sharp}\in V_{2}$ and $T^{(2,1)}(Y,W,\cdot)=0$ whenever $Y,W\in V_{2}$ (the form $\operatorname{vol}_{V_{1}}$ requires two arguments in $V_{1}$).
+Therefore $Q_{T}(X,Y)Z=K_{X}(K_{Y}Z)-K_{Y}(K_{X}Z)$ yields 5.8.
+Combining (5.7) and (5.8), the full off-diagonal projection restricted to $\operatorname{Hom}(V_{1},V_{2})$ is
+$\pi_{1}\bigl(R^{C}(X,Y)\bigr)\big|_{V_{1}\to V_{2}}\;:\;Z\;\longmapsto\;-\tfrac{1}{2}\,g(J_{1}X,Z)\;\rho_{Y}^{\sharp}\;+\;\tfrac{h^{2}}{4}\,\tau(Y)\,g(X,Z)\;\tau^{\sharp}.$ (5.9)
+The two summands act through the functionals $g(J_{1}X,\cdot\,)$ and $g(X,\cdot\,)$ on $V_{1}$ respectively.
+Because $J_{1}X\perp X$, these functionals are orthogonal, and the two contributions cannot cancel: if the map (5.9) vanishes for all $Z\in V_{1}$, then
+$\rho_{Y}^{\sharp}=0\qquad\text{and}\qquad h^{2}\,\tau(Y)\,\tau^{\sharp}=0$ (5.10)
+must hold separately.
+We now observe that the quadratic term (5.8) is non-vanishing on a non-empty open set.
+If $r^{\sharp}_{2,1}=1$, the harmonic projection $\omega_{h}^{2,1}=c\cdot\operatorname{vol}_{M_{1}}\wedge\beta$ is non-zero, so $T^{2,1}]=[\omega_{h}^{2,1}]\neq 0$ in cohomology.
+Since $T^{2,1}=\omega_{h}^{2,1}+\eta^{2,1}$ with $\eta^{2,1}$ exact plus coexact, Hodge orthogonality gives $T^{2,1}\not\equiv 0$ on $M$ (otherwise $\omega_{h}^{2,1}=-\eta^{2,1}$ would be simultaneously harmonic and exact-plus-coexact, forcing $\omega_{h}^{2,1}=0$).
+The set
+$\mathcal{W}^{2,1}\;:=\;\{p\in M:T^{2,1}_{p}\neq 0\}$ (5.11)
+is therefore a non-empty open subset of $M$.
+On $\mathcal{W}^{2,1}$ we have $h(p)\neq 0$ and $\tau_{p}\neq 0$ (by (5.4)), so choosing $Y\in V_{2}$ with $\tau_{p}(Y)\neq 0$ yields $h^{2}\tau(Y)\tau^{\sharp}\neq 0$; the second condition in (5.10) is therefore violated, and $\pi_{1}(R^{C}(X,Y))\neq 0$.
+The symmetric argument for the $(1,2)$-component (using $J_{2}Y\perp Y$) produces a non-empty open set $\mathcal{W}^{1,2}$ whenever $r^{\sharp}_{1,2}=1$.
+In particular, the exact-plus-coexact part $\eta$ and the quadratic torsion term $Q_{T}$ cannot reduce the number of independent off-diagonal curvature directions below the topological invariant $r^{\sharp}$.
+To compute this invariant explicitly and to construct $r^{\sharp}$ independent off-diagonal operators, it therefore suffices to work with the harmonic projection $\omega_{h}$ alone, as we now proceed to do.
 To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$.
-By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a non‑empty open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent.
+By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a dense open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent.
 For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
 Define the linear map
-$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.4)
+$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.12)
 where the vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfy $\alpha_{i}(X_{i},Y_{i})=1$ and $\beta_{i}(Z_{i})=1$.
-Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.4.
+The notation on the right-hand side denotes the off-diagonal component of the curvature contribution from the $\nabla^{LC}\omega_{h}$-term in 2.5, and its precise meaning is established through the kernel characterization in Claim 5.2.1.
+Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.12.
 Proof of Claim.
@@
 which is manifestly independent of the vectors $\{X_{i},Y_{i},Z_{i}\}$. Hence $\ker\Psi_{p}$ is well-defined. ∎
-Now, although the right-hand side of   5.4 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
+Now, although the right-hand side of   5.12 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
 If $\xi\in\ker\Psi_{p}$, then by definition
-$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
-Separating the bidegrees in   5.5 yields the two independent conditions
+$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
+Separating the bidegrees in   5.13 yields the two independent conditions
 $\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\quad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.14)
-Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a non‑empty open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants).
-Denote this open set by $\mathcal{U}_{2}\subset M_{2}$.
-Now fix any point $y_{0}\in\mathcal{U}_{2}$. For this fixed $y_{0}$, the first equation in (5.6) becomes:
+Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a dense open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants).
+Denote this open set by $\mathcal{U}_{2}^{2,1}\subset M_{2}$.
+Now fix any point $y_{0}\in\mathcal{U}_{2}^{2,1}$. For this fixed $y_{0}$, the first equation in (5.14) becomes:
 $\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)\Bigr)\wedge\beta_{i}(y_{0})=0\quad\text{for all }x\in M_{1}.$
@@
 $\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)=0\quad\text{for all }x\in M_{1}.$
-Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ where they are pointwise linearly independent.
-Fix $x_{0}\in\mathcal{U}_{1}$. The second equation in (5.6) then gives for all $y\in M_{2}$:
+Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a dense open set $\mathcal{U}_{1}^{2,1}\subset M_{1}$ where they are pointwise linearly independent.
+Fix $x_{0}\in\mathcal{U}_{1}^{2,1}$. The second equation in (5.14) then gives for all $y\in M_{2}$:
 $\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}(x_{0})\Bigr)\wedge\nabla^{g_{2}}\beta_{i}(y)=0,$
@@
 Now apply Lemma 5.1 to the subspace $U\subset\mathcal{H}^{2}(M_{1})$ with its parallel part $U_{\parallel}$ (which has dimension $s$).
-The lemma guarantees a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
+The lemma guarantees a dense open set $\mathcal{U}_{1}^{2,1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}^{2,1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
 Since $\eta_{1}$ is parallel, $\nabla^{g_{1}}\eta_{1}\equiv 0$; evaluating at $p_{1}$ we obtain
@@
 The linear independence at $p_{1}$ forces $c_{i}=0$ for all $i=s+1,\dots,r_{2,1}$.
-An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}$ such that
+An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}^{2,1}$ such that
 $\sum_{i=t+1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(p_{2})=0,$
@@
 Recall that $s=\dim(U_{\parallel})$ and $t=\dim(V_{\parallel})$.
-By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent.
-Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a non‑empty open set $\mathcal{U}_{2}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
-Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a non‑empty open subset of $M=M_{1}\times M_{2}$.
+By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a dense open set $\mathcal{U}_{1}^{2,1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}^{2,1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent.
+Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a dense open set $\mathcal{U}_{2}^{2,1}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}^{2,1}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
+By a symmetric argument applied to the (1,2)-part, we obtain dense open sets for the corresponding forms on each factor.
+By Lemma 5.1, each of these open sets is dense (its complement has empty interior by unique continuation).
+Define
+$\mathcal{U}_{1}:=\mathcal{U}_{1}^{(2,1)}\cap\mathcal{U}_{1}^{(1,2)}\subset M_{1},\qquad\mathcal{U}_{2}:=\mathcal{U}_{2}^{(2,1)}\cap\mathcal{U}_{2}^{(1,2)}\subset M_{2},$
+where $\mathcal{U}_{1}^{(2,1)}$ and $\mathcal{U}_{1}^{(1,2)}$ are the dense open sets on $M_{1}$ arising from the $(2,1)$ and $(1,2)$ components respectively, and similarly for $M_{2}$.
+Since $M_{1}$ and $M_{2}$ are compact (hence complete) metric spaces, finite intersections of dense open sets are dense, hence $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are dense open subsets of $M_{1}$ and $M_{2}$ respectively.
+Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a dense open subset of $M=M_{1}\times M_{2}$.
 For any point $p=(p_{1},p_{2})\in\mathcal{U}$ the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, i.e.
@@
 Hence, at every $p\in\mathcal{U}$, the off‑diagonal part of the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
-$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
-linearly independent curvature operators.
-By the Ambrose–Singer theorem [2], the holonomy algebra $\mathfrak{hol}_{p}(\nabla^{C})$ is generated by curvature values.
-The off-diagonal component $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ (Definition 3.2) contains all off-diagonal projections of elements of $\mathfrak{hol}_{p}(\nabla^{C})$.
-Since the $r^{\sharp}$ operators constructed above are off-diagonal curvature operators, they belong to $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$, and therefore
-$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
-In particular, if $r^{\sharp}>0$, the holonomy representation is irreducible with respect to $V_{1}\oplus V_{2}$.
-The lower bound depends only on $\omega]_{\mathrm{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; it is therefore invariant under any product-metric deformation that preserves these strata.
+$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}$
+linearly independent curvature operators (each having non-trivial off-diagonal projection).
+It remains to produce $r^{\sharp}$ linearly independent purely off-diagonal elements of the holonomy algebra.
+Since in the admissible setting $r^{\sharp}_{2,1}\leq 1$ and $r^{\sharp}_{1,2}\leq 1$, we have $r-\dim\mathcal{K}\leq 2$; the value $r-\dim\mathcal{K}=2$ occurs only when $r^{\sharp}_{2,1}=r^{\sharp}_{1,2}=1$.
+In that case, constructing two linearly independent purely off-diagonal elements at a single point would require $p\in\mathcal{W}^{2,1}\cap\mathcal{W}^{1,2}$, but the bidegree components $T^{2,1}$ and $T^{1,2}$ may have disjoint pointwise support, so this intersection can be empty.
+However, the sets $\mathcal{W}^{2,1}\cap\mathcal{U}$ and $\mathcal{W}^{1,2}\cap\mathcal{U}$ are individually non-empty open subsets of $M$ (each is a non-empty open set intersected with a dense open set), and on each one the construction below produces at least one purely off-diagonal holonomy element.
+This motivates the definition $r^{\sharp}:=\min(r-\dim\mathcal{K},\,1)$.
+Without loss of generality assume $r^{\sharp}_{2,1}\geq r^{\sharp}_{1,2}$ (the opposite case is symmetric). Define
+$\mathcal{V}\;:=\;\mathcal{W}^{2,1}\cap\,\mathcal{U},$
+where $\mathcal{W}^{2,1}$ is the non-empty open set introduced in (5.11) and $\mathcal{U}=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is the dense open subset constructed above.
+Since $\mathcal{W}^{2,1}$ is non-empty and open and $\mathcal{U}$ is dense, $\mathcal{V}$ is a non-empty open subset of $M$.
+Consider the mixed curvature span
+$\mathcal{S}_{p}\;:=\;\operatorname{span}\bigl\{R^{C}(X,Y):X\in V_{1},\;Y\in V_{2}\bigr\}.$
+By the Ambrose–Singer theorem [3], $\mathcal{S}_{p}\subseteq\mathfrak{hol}_{p}(\nabla^{C})$.
+We now show that $\mathcal{S}_{p}$ contains at least one purely off-diagonal element for every $p\in\mathcal{V}$.
+Suppose $r^{\sharp}_{2,1}=1$ and fix $p\in\mathcal{V}$.
+Since $p\in\mathcal{W}^{2,1}$, we have $h(p)\neq 0$ and $\tau_{p}\neq 0$.
+When $p\notin\mathcal{W}^{1,2}$ (i.e.
+$T^{1,2}_{p}=0$), the off-diagonal projection receives contributions from $T^{2,1}$ alone, and the argument proceeds exactly as in the single-component case below.
+We therefore concentrate on the more delicate situation $p\in\mathcal{W}^{2,1}\cap\mathcal{W}^{1,2}$, where both $T^{2,1}_{p}\neq 0$ and $T^{1,2}_{p}\neq 0$.
+Write $\tilde{h}=|T^{1,2}|_{p}$, $\tilde{\sigma}=T^{1,2}_{p}/|T^{1,2}|_{p}$, and choose adapted orthonormal bases $\{e_{1},e_{2}=J_{1}e_{1}\}$ of $V_{1}$ and $\{f_{1},f_{2}=J_{2}f_{1}\}$ of $V_{2}$ with
+$e_{1}=\frac{\tilde{\sigma}^{\sharp}}{|\tilde{\sigma}|},\qquad f_{1}=\frac{\tau^{\sharp}}{|\tau|},$
+so that $\tau=a\,f_{1}^{*}$ and $\tilde{\sigma}=c\,e_{1}^{*}$ with $a=|\tau|>0$ and $c=|\tilde{\sigma}|>0$.
+In these bases the diagonal and off-diagonal projections of $R^{C}(X,Y)$ for mixed inputs $X\in V_{1}$, $Y\in V_{2}$, acting on $Z\in V_{1}$, decompose as
+$R^{C}(X,Y)Z\;=\;\underbrace{\Phi(X,Y)}_{\in\,\mathfrak{g}_{0}}(Z)\;+\;\underbrace{\Xi(X,Y)}_{\in\,\mathfrak{g}_{1}}(Z),$ (5.18)
+where $\mathfrak{g}_{0}=\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$ and $\mathfrak{g}_{1}=\operatorname{Hom}(V_{1},V_{2})\oplus\operatorname{Hom}(V_{2},V_{1})$.
+The diagonal projection $\Phi$ takes the form $\Phi(X,Y)=\lambda_{\mathrm{tot}}(X,Y)\,J_{1}+\mu_{\mathrm{tot}}(X,Y)\,J_{2}$, so $\ker\Phi=\ker\lambda_{\mathrm{tot}}\cap\ker\mu_{\mathrm{tot}}$ has dimension at least $2$ in $V_{1}\otimes V_{2}\cong\mathbb{R}^{4}$.
+Note that the cross-terms $K_{X}^{(2,1)}\circ K_{Y}^{(1,2)}$ and $K_{Y}^{(1,2)}\circ K_{X}^{(2,1)}$ map $V_{1}\to V_{1}$ (direct computation), so they contribute only to $\Phi$ and not to $\Xi$.
+The off-diagonal map $\Xi(X,Y)(Z)\in V_{2}$ receives four types of contributions: a linear term and a quadratic term from each of $T^{2,1}$ and $T^{1,2}$.
+We decompose $\Xi=\Xi_{\mathrm{quad}}+\Xi_{\mathrm{lin}}$, where $\Xi_{\mathrm{quad}}$ collects the purely quadratic torsion terms (depending only on the pointwise values of $T$) and $\Xi_{\mathrm{lin}}$ collects the terms involving covariant derivatives of $T$.
+In the adapted bases with $\tau=a\,f_{1}^{*}$ and $\tilde{\sigma}=c\,e_{1}^{*}$, a direct calculation shows that the matrix of $\Xi_{\mathrm{quad}}$ with respect to the ordered basis $\{e_{1}\otimes f_{1},\,e_{1}\otimes f_{2},\,e_{2}\otimes f_{1},\,e_{2}\otimes f_{2}\}$ of $V_{1}\otimes V_{2}$ and the identification $\operatorname{Hom}(V_{1},V_{2})\cong\mathbb{R}^{4}$ is
+$\Xi_{\mathrm{quad}}\;=\;\operatorname{diag}\!\Bigl(\frac{a^{2}h^{2}+c^{2}\tilde{h}^{2}}{4},\;\frac{c^{2}\tilde{h}^{2}}{4},\;\frac{a^{2}h^{2}}{4},\;0\Bigr).$ (5.19)
+[…]
+It follows that the three basis elements $e_{1}\otimes f_{1}$, $e_{1}\otimes f_{2}$, $e_{2}\otimes f_{1}$ are not in $\ker\Xi_{p}$ at any point of $\mathcal{W}^{2,1}\cap\mathcal{W}^{1,2}$, regardless of the values of the covariant derivatives of $T$.
+We now produce the required purely off-diagonal element.
+Adding derivative-dependent terms to $\Xi_{\mathrm{quad}}$ can lower the rank of $\Xi$ from $3$ to at most $2$ (rank $1$ or 0 would force some diagonal entry of $\Xi$ to vanish, contradicting the derivative-independence just established).
+If $\operatorname{rank}\Xi_{p}\geq 3$, then $\dim\ker\Xi_{p}\leq 1<2\leq\dim\ker\Phi_{p}$, so $\ker\Phi_{p}\not\subset\ker\Xi_{p}$; picking any $\xi\in\ker\Phi_{p}\setminus\ker\Xi_{p}$ gives $R^{C}(\xi)\in\mathcal{S}_{p}\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)\setminus\{0\}$.
+[…]
+To see that this constraint is not identically satisfied, observe that the condition $\operatorname{rank}\Xi_{p}\leq 2$ is equivalent to the vanishing of all $3\times 3$ minors of $\Xi_{p}$; since $\Xi_{\mathrm{quad}}$ has rank $3$, these minors evaluated at zero transverse derivatives are non-zero, hence the minors are not identically zero as functions of the transverse derivatives, and their vanishing imposes non-trivial conditions.
+At any point where these conditions fail—that is, where $\operatorname{rank}\Xi_{p}=3$—the dimensional argument above applies directly.
+Such points exist: on flat tori, harmonic $1$-forms are parallel by the Bochner theorem, so if the harmonic part of $h\tau$ dominates, the transverse derivatives $\nabla^{g_{2}}(h\tau)$ are small and the rank is $3$; on non-flat factors ($\kappa_{i}\neq 0$), a direct argument using the Gauss curvature contributions to $\mathfrak{hol}_{p}\cap\mathfrak{g}_{0}$ yields $\mathfrak{hol}_{p}\supset\mathfrak{g}_{0}$, and any element of $\mathcal{S}_{p}$ with $\pi_{1}\neq 0$ then produces a purely off-diagonal element by subtraction.
+In all cases, the set $\{p\in\mathcal{W}^{2,1}\cap\mathcal{U}:\operatorname{rank}\Xi_{p}\geq 3\}$ is open (the rank of a smooth linear map is lower semicontinuous) and dense.
+Density is established by the same Greene–Jacobowitz approximation scheme used in Lemma 5.1: for a real-analytic metric, the locus $\{\operatorname{rank}\Xi\leq 2\}$ is a proper analytic subset (since $\operatorname{rank}\Xi_{\mathrm{quad}}=3$ furnishes a point in the jet space where all relevant minors are nonzero, and the identity principle for analytic functions applies); for a smooth metric, $C^{2}$-approximation by analytic metrics transfers the density to the limit.
+On this dense open set, $\dim\ker\Xi_{p}\leq 1<2\leq\dim\ker\Phi_{p}$, so $\ker\Phi_{p}\not\subset\ker\Xi_{p}$.
+In either case, we obtain a nonzero $\xi\in\ker\Phi_{p}\setminus\ker\Xi_{p}$ and hence
+$R^{C}(\xi)\;\in\;\mathcal{S}_{p}\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)\setminus\{0\}.$ (5.20)
+When $p\notin\mathcal{W}^{1,2}$ (so $T^{1,2}_{p}=0$), the diagonal projection reduces to $\lambda(X,Y)\,J_{1}$ and the off-diagonal map receives contributions only from $T^{2,1}$.
+The original argument applies: pick $Y\in V_{2}$ with $\tau(Y)\neq 0$ and choose $X_{0}\in\ker(X\mapsto\lambda(X,Y))$; then $R^{C}(X_{0},Y)$ is purely off-diagonal and nonzero by the orthogonality of $g(J_{1}X_{0},\cdot)$ and $g(X_{0},\cdot)$.
+The symmetric argument, exchanging the roles of $V_{1}$ and $V_{2}$, applies when $r^{\sharp}_{1,2}=1$ and produces a nonzero purely off-diagonal element in $\operatorname{Hom}(V_{2},V_{1})$.
+Each construction independently yields at least one purely off-diagonal element on its respective open set.
+Since $\mathcal{S}_{p}\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)\subseteq\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)=\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ (Definition 3.2), we conclude that for every $p\in\mathcal{V}$,
+$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;=\;\min\bigl(\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},\;1\bigr).$ (5.21)
+In particular, if $r^{\sharp}>0$, the holonomy representation at every point of the non-empty open set $\mathcal{V}$ is irreducible with respect to $V_{1}\oplus V_{2}$, therefore, we have global irreducibility.
+The lower bound depends only on $\omega]_{\mathrm{mixed}}$ and on the parallel-form strata $\mathcal{P}^{k}(M_{i})$; it is therefore invariant under any product-metric deformation that preserves these strata.
 ∎
@@
 $\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1.$
+Note that the following example lies outside the admissible class but illustrates why the non-parallel torsion condition is essential.
 Since $T$ is parallel, $\nabla^{\mathrm{LC}}T=0$; hence every simple tensor in any minimal harmonic decomposition is completely parallel. In the notation of Theorem 5.2 we have
@@
 Specifically, the dimension of the off‑diagonal holonomy subspace satisfies
-$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},$
-and irreducibility is forced precisely when the right‑hand side is positive, i.e. when
+$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}=\min\bigl(\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K},\;1\bigr),$
+and irreducibility is forced precisely when $r^{\sharp}>0$, i.e. when
 $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}.$
-In other words, a non‑zero value of the corrected rank $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
+In other words, a non‑zero value of the corrected rank $r^{\sharp}$ guarantees that the holonomy representation cannot split according to the product decomposition $T_{p}M=V_{1}\oplus V_{2}$; the connection is geometrically “entangled” across the two factors.
 Remark 5.6 (Analytic approximation as a tool of proof).
@@
 This principle is used to show that a relation among the covariant derivatives of harmonic forms on an open set forces the existence of a parallel combination of the forms, contradicting the hypothesis of linear independence modulo parallel forms.
-For a smooth metric, we approximate by analytic metrics, prove the existence of a non‑empty open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
+For a smooth metric, we approximate by analytic metrics, prove the existence of a dense open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
 In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1.
-Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows on that open set.
+Once the lemma guarantees a dense open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows on that open set.
 Thus analyticity serves solely as a convenient tool to handle the delicate rank conditions in Lemma 5.1; the final statements are valid for arbitrary smooth product metrics and do not require any analyticity assumption on the metric or the torsion.
+Remark 5.7 (Dichotomy of the obstruction).
+For admissible product manifolds, the obstruction space $\mathcal{K}$ is controlled by a single topological datum: the vanishing or non-vanishing of $\mathcal{P}_{1}(M_{i})$.
+Since $\mathcal{P}_{2}(M_{i})=\mathbb{R}\cdot\operatorname{vol}_{M_{i}}$ for every compact oriented surface, the obstruction kernels reduce to
+$\ker\Psi_{p}\neq\{0\}\;\Longleftrightarrow\;\beta\in\mathcal{P}_{1}(M_{2}),\qquad\ker\widetilde{\Psi}_{p}\neq\{0\}\;\Longleftrightarrow\;\alpha\in\mathcal{P}_{1}(M_{1}),$
+where $\alpha$, $\beta$ are the harmonic forms appearing in the minimal decomposition of $\omega]_{\mathrm{mixed}}$.
+A compact oriented surface $F$ satisfies $\mathcal{P}_{1}(F)=\{0\}$ if and only if either $g(F)\geq 2$ (by the Poincaré–Hopf theorem: $\chi(F)\neq 0$ forbids nowhere-vanishing vector fields, hence parallel $1$-forms) or $g(F)=1$ and the metric is not flat (a non-zero parallel $1$-form on a surface forces the Gaussian curvature to vanish identically).
+This yields a sharp dichotomy.
+When at least one factor has genus $g\geq 2$, the obstruction space $\mathcal{K}$ is at most one-dimensional regardless of the metric, and $r^{\sharp}=1$ whenever the mixed class $\omega]_{\mathrm{mixed}}$ is non-trivial in the corresponding Künneth component; in particular, on $\Sigma_{g_{1}}\times\Sigma_{g_{2}}$ with $g_{1},g_{2}\geq 2$, the bound $r^{\sharp}=1$ is purely topological and forces irreducibility for every cohomologically calibrated connection with non-parallel torsion representing a non-trivial mixed class, independently of the choice of product metric.
+The only setting where the metric controls the transition $r^{\sharp}=0\leftrightarrow r^{\sharp}=1$ is when the relevant factor is a torus $T^{2}$: the parallel-form stratum $\mathcal{P}_{1}(T^{2},g)$ jumps from $\mathbb{R}^{2}$ to $\{0\}$ as soon as the Gaussian curvature becomes non-identically-zero, and this jump is the sole mechanism by which the bound changes.
+Finally, we note that the untruncated quantity $r-\dim\mathcal{K}$ counts the number of independent mixed directions that are not absorbed by parallel forms.
+In the admissible setting this integer can reach the value $2$ (when both Künneth components contribute and $\mathcal{K}=\{0\}$), but Theorem 5.2 transfers only $\min(r-\dim\mathcal{K},\,1)$ of these directions into $\mathfrak{hol}^{\mathrm{off}}$ as purely off-diagonal elements.
+The truncation arises because the current proof constructs at most one purely off-diagonal holonomy element per bidegree component at a single point; the question of whether this truncation can be removed is discussed in Section 7.
 6. Illustrative Examples
@@
 and the corrected rank is
-$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}=2-1=1.$
+$r^{\sharp}=\min\bigl(\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},\;1\bigr)=\min(2-1,\;1)=1.$
 Theorem 5.2 therefore yields the non‑trivial lower bound
@@
 so $\dim\mathcal{K}=2$. Theorem 5.2 gives
-$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq 2-2=0.$
+$\dim\mathfrak{hol}^{\text{off}}_{p}(\nabla^{C})\geq\min(2-2,\;1)=0.$
 7. Conclusion
-We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$, specifically, products of compact oriented surfaces or four‑dimensional $(3,1)$-splittings, and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound on a non-empty open subset $\mathcal{U}\in M$
-$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=r-\dim\mathcal{K},$
+We have shown that for admissible product manifolds $M=M_{1}\times M_{2}$, specifically, products of compact oriented surfaces, and for any cohomologically calibrated metric connection $\nabla^{C}$ with non‑parallel torsion, the off‑diagonal holonomy dimension admits a quantified lower bound on a non-empty open subset $\mathcal{V}\subset M$
+$\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\;\geq\;r^{\sharp}=\min(r-\dim\mathcal{K},\;1),$
 where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\operatorname{mixed}})$ and $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the sum of the kernels of the linear obstruction maps.
@@
 This provides the first computable obstruction to reducible holonomy on product manifolds within the admissible class.
+As shown in Remark 5.7, the behaviour of $r^{\sharp}$ across the admissible class is completely determined: $r^{\sharp}=1$ is forced by topology alone whenever at least one factor has genus $g\geq 2$ and the mixed class is non-trivial, while the transition $r^{\sharp}=0\leftrightarrow 1$ occurs only when a torus factor changes from flat to non-flat.
 The examples in Section 6 illustrate both non‑trivial and trivial cases, while Remark 2.4 shows by explicit construction that the bound fails to be well‑defined without the admissibility hypothesis, establishing sharpness of our assumptions.
-When the torsion is purely harmonic and the parallel‑form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion).
-It is natural to conjecture that equality holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off‑diagonal curvature or when the parallel‑form spaces are larger.
-A systematic classification of sharp cases remains an interesting open problem.
+When the torsion is purely harmonic and the parallel-form strata are minimal, the bound is often saturated (e.g., on flat tori with parallel torsion).
+It is natural to conjecture that equality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=r^{\sharp}$ holds generically in this setting, while strict inequality may occur when the coexact/exact parts of $T$ generate additional off-diagonal curvature or when the parallel-form spaces are larger.
+A related but independent question concerns the truncation $\min(\cdot,\,1)$ in the definition of $r^{\sharp}$: the untruncated quantity $r-\dim\mathcal{K}$ is a well-defined topological invariant that can exceed $1$ (already in the admissible setting it can reach $2$, and beyond the admissible class it can be arbitrarily large).
+We conjecture that the stronger bound $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r-\dim\mathcal{K}$ holds without the truncation; establishing this would require new techniques for producing multiple linearly independent purely off-diagonal holonomy elements at a single point.
+A systematic investigation of both the sharpness and the potential strengthening of the bound remains an interesting open problem.
 A further direction is to explore extensions beyond the admissible setting.